### v1

# Bayesian Inference for Consistent Predictions in Overparameterized Nonlinear Regression

Abstract.
The remarkable generalization performance of overparameterized models has challenged the conventional wisdom of statistical learning theory. While recent theoretical studies have shed light on this behavior in linear models or nonlinear classifiers, a comprehensive understanding of overparameterization in nonlinear regression models remains lacking. This paper explores the predictive properties of overparameterized nonlinear regression within the Bayesian framework, extending the methodology of adaptive prior based on the intrinsic spectral structure of the data. We establish posterior contraction for single-neuron models with Lipschitz continuous activation functions and for generalized linear models, demonstrating that our approach achieves consistent predictions in the overparameterized regime. Moreover, our Bayesian framework allows for uncertainty estimation of the predictions. The proposed method is validated through numerical simulations and a real data application, showcasing its ability to achieve accurate predictions and reliable uncertainty estimates. Our work advances the theoretical understanding of the blessing of overparameterization and offers a principled Bayesian approach for prediction in large nonlinear models.
1. Introduction
In recent years, the field of machine learning has witnessed a paradigm shift in understanding the generalization performance. Traditional statistical theory emphasized the importance of controlling model complexity and avoiding overfitting by balancing the bias-variance tradeoff [55]. However, the emergence of large-scale deep neural networks, which often have far more parameters than the number of training data, has challenged this conventional wisdom. These models have demonstrated the ability to achieve high accuracy on test data while completely interpolating the training data. This discrepancy between these empirical observations and traditional learning theory has necessitated a new explanation [61].
A groundbreaking study by [5] provided a theoretical foundation in the overparameterized linear regression models, showing the non-asymptotic analysis of the generalization error of a minimum norm interpolator. This discovery has reshaped our understanding of generalization performance in large statistical models, suggesting that overparameterization does not necessarily lead to poor prediction, even in simple models. Instead, the key to good performance lies in the model’s expressiveness and its ability to capture the essential structure of the data. Indeed, subsequent theoretical studies have demonstrated that overparameterization leads to accurate predictions in various statistical models, shedding light on the behavior of modern machine learning models.
While the theoretical understanding of the goodness of overparameterization has advanced significantly in linear estimators, such as linear regression [5, 52] and kernel methods [31, 33], the analysis of nonlinear estimators remains limited. Although some studies have investigated this issue in nonlinear neural networks for classification problems [13, 26], a comprehensive understanding of why and under what conditions overparameterized nonlinear regression can achieve good generalization performance is still lacking. Hence, addressing this question is crucial for reinforcing the theoretical foundations of modern machine learning.
In this paper, we explore the predictive properties of overparameterized nonlinear regression models within the Bayesian framework. Bayesian inference combines the prior distribution of parameters with a likelihood function to obtain the posterior distribution. [59] proposed a methodology that incorporates the structure of the data into the prior distribution while capturing the linear model structure as the likelihood, resulting in accurate predictions even in non-sparse high-dimensional models. The key idea is to adaptively adjust the prior variance based on the intrinsic structure (spectra) of the data, which is found to be critical in prediction error convergence in the sense of posterior contraction. We extend this to nonlinear regression models, specifically single neuron models [60, 47] and generalized linear models (GLMs) [37], and theoretically analyze the prediction error of the resulting posterior distribution. The main contributions of this study are summarized as follows:
  • For generalized linear models based on one-parameter exponential families, including logistic and Poisson regressions, we establish conditions under which the posterior distribution converges to the true parameter in terms of the predictive risk and provide an upper bound on the convergence rate (Theorem 1). This finding highlights the applicability of our approach to the regression of discrete response variables, such as those encountered in classification problems.
  • We show that the proposed method generalizes well in single-neuron models with a Lipschitz continuous nonlinear activation function and Gaussian noise (Theorem 2). This result suggests that, even in more general models involving nonlinearity, overparameterization allows us to predict benignly under appropriate prior distributions.
  • Our Bayesian approach not only provides a theoretical guarantee of the generalization performance of the overparametrized models but also estimates the uncertainty in the predictions. By sampling from the posterior distribution, we can construct a confidence level for each prediction, which is essential in decision-making processes.
1.1. Related work
Bayesian GLM in high-dimension has been studied intensively. The first paper to study high-dimensional Bayesian GLM [18] achieved posterior concentration for the parameter without assuming any sparsity structure. However, it imposed the restricted growth condition on the parameter dimension, and this underparameterized setting differs from the present work. Subsequently, several studies have explored the GLM setting where the number of parameters may exceed the sample size [24, 34, 23]. These works have demonstrated the asymptotic properties of the posterior distribution, but they assume sparsity, implying that the parameters are inherently low-dimensional. The properties of Bayesian variable selection in GLMs have also been investigated in similar domains [8, 9, 42, 30, 22]. However, the assumption of parameter sparsity can be stringent, and the current study is conducted in a setting that does not rely on this assumption, thus expanding the scope of research in this area. The methods and theory of Bayesian inference in high-dimensional models are summarized in [2].
Numerous theoretical results suggest that overparameterization leads to good generalization performance. [5] demonstrated that the excess risk converges to zero in overparameterized linear models. This theory was widely extended in other linear estimators, such as regularized regression [52, 29, 58, 36, 11], kernel regression [31, 32], linear neural networks [10, 50], and regression for dependent data [40, 53]. [13, 6, 26] have ventured beyond these linear frameworks to investigate the benign nature of overparameterization, but their focus is limited to classification problems. [47] tackled nonlinear regression and demonstrated the absence of benign overfitting, but their assumptions differ from the present work mainly in terms of distributional assumptions on the covariate.
1.2. Organization
The remainder of this paper is structured as follows. In Section 2, we introduce the problem setting of one-parameter generalized linear models under the overparameterization assumption and provide a detailed description of our proposed methodology. Section 3 presents the theoretical results on posterior contraction for GLMs. Section 4 extends the results to a two-parameter case. We then conduct simulation studies in Section 5 to validate our theoretical findings and assess the uncertainty estimation. Section 6 showcases a real data application, comparing our approach with baseline methods in terms of classification performance. In Section 7, we conclude this paper by discussing the contribution and limitations of this work. The appendix provides the proof and discussion of the main theorems and the details of the computational algorithms used in our analysis.
1.3. Notation
For $q\in\mathbb{N}$ and any vector $\bm{v}$, $\|\bm{v}\|_{q}$ denotes the $\ell_{q}$-norm of $\bm{v}$. For any vector $\bm{v}$ and symmetric matrix of the same dimension $A$, we define $\|\bm{v}\|_{A}=\sqrt{\bm{v}^{\top}A\bm{v}}$. For sequences $\{a_{n}\}_{n\in\mathbb{N}}$ and $\{b_{n}\}_{n\in\mathbb{N}}$, $a_{n}\gtrsim b_{n}$ denotes the existence of a constant $c>0$ such that $a_{n}\geq cb_{n}$ for all $n\geq\overline{n}$ with some finite $\overline{n}\in\mathbb{N}$, and $a_{n}\lesssim b_{n}$ denotes the opposite. $a_{n}=o(b_{n})$ indicates that $a_{n}/b_{n}\to 0$ as $n\to\infty$ whereas $a_{n}=\omega(b_{n})$ does $a_{n}/b_{n}\to\infty$ as $n\to\infty$. $a_{n}\gtrapprox b_{n}$ denotes that $a_{n}\gtrsim b_{n}(\log n)^{c}$ holds for every $c>0$, and $a_{n}\lessapprox b_{n}$ denotes the opposite. For natural numbers $a,b\in\mathbb{N}$ with $a<b$, we define $[a:b]:=\{a,a+1,a+2,...,b\}$. For any square matrix $A$, let $A^{-1}$ be the inverse matrix of $A$, $A^{\dagger}$ be the Moore–Penrose pseudoinverse matrix, and $\mathrm{tr}(A)$ be the sum of the diagonals of $A$. For a pseudo-metric space $(S,d)$ and a positive value $\delta$, $\mathcal{N}(\delta,S,d)$ denotes the covering number, that is, the minimum number of balls that cover $S$ with a radius $\delta$ in terms of $d$. For two probability measures $\Pi_{1}$ and $\Pi_{2}$ on the same measurable space $(\mathcal{X},\mathcal{B})$, the total variation distance is given by $\|\Pi_{1}-\Pi_{2}\|_{\mathrm{TV}}:=\sup_{B\in\mathcal{B}}|\Pi_{1}(B)-\Pi_{2}(B)|$, and the Kullback–Leibler (KL) divergence and the KL variation (also referred to as the centered second-order KL divergence) between them are respectively defined as $K(\Pi_{1},\Pi_{2})=\int_{\mathcal{X}}\log(\mathrm{d}\Pi_{1}/\mathrm{d}\Pi_{2})\mathrm{d}\Pi_{1}$ and $V(\Pi_{1},\Pi_{2})=\int_{\mathcal{X}}\{\log(\mathrm{d}\Pi_{1}/\mathrm{d}\Pi_{2})-K(\Pi_{1},\Pi_{2})\}^{2}\mathrm{d}\Pi_{1}$ if $\Pi_{1}$ is absolutely continuous with respect to $\Pi_{2}$; otherwise, both are $\infty$.∞ . The Hellinger distance between two probability distributions $\Pi_{1}$ and $\Pi_{2}$ is defined as $H(\Pi_{1},\Pi_{2})=\sqrt{1-\int_{\mathcal{X}}\sqrt{\mathrm{d}\Pi_{1}\mathrm{d}\Pi_{2}}}$.
2. One-parameter Generalized Linear Regression Model
This section presents a comprehensive overview of the one-parameter generalized linear regression model, including the problem setting (§2.1) and the composition of the predictor (§2.2).
2.1. Setting
Consider a pair $(\bm{X},Y)$, where $\bm{X}$ is an $\mathbb{R}^{p}$-valued centered random covariate and $Y$ is an $\mathbb{R}$-valued univariate response. The GLM is defined as follows:
$\mathbb{E}[Y\mid\bm{X}=\bm{x}]=g(\bm{x}^{\top}\bm{\beta}^{*}),$ (1)
where $\bm{\beta}^{*}\in\mathbb{R}^{p}$ is the true unknown $p$-dimensional regression coefficient, and $g:\mathbb{R}\to\mathbb{R}$ is a known Lipschitz continuous function. In this setting, the response variable $Y$ follows a general distribution belonging to the one-parameter exponential family, with a density function given by:
$L(y;\mu_{x,\bm{\beta}})=\exp\{a(\mu_{x,\bm{\beta}})y+b(\mu_{x,\bm{\beta}})+c(y)\},$ (2)
where $\mu_{x,\bm{\beta}}\in\mathbb{R}$ is the parameter, $a:\mathbb{R}\to\mathbb{R}$ and $b:\mathbb{R}\to\mathbb{R}$ are known continuously differentiable functions, with $a$ having a nonzero derivative. The function $c:\mathbb{R}\to\mathbb{R}$ is chosen such that $L(y;\mu_{x,\bm{\beta}})$ is a proper probability density.
This formulation encompasses several well-known models: Normal linear models: $Y\mid\bm{X}$ follows a $g(\bm{x}^{\top}\bm{\beta}^{*})$-mean Gaussian distribution with fixed variance, and $g$ is the identity function; Logistic regression model: $Y\mid\bm{X}$ follows a Bernoulli distribution with parameter $g(\bm{x}^{\top}\bm{\beta}^{*})$, and $g$ is the logistic function; Poisson regression model: $Y\mid\bm{X}$ follows a Poisson distribution with parameter $g(\bm{x}^{\top}\bm{\beta}^{*})$, and $g$ is the exponential function. Note that we do not restrict the link function $g$ to canonical, providing flexibility in modeling the relationship between the linear predictor and the response variable.
This study focuses on a non-sparse overparameterization setting, where $n$ independent and identically distributed copies of $(\bm{X},Y)$ are generated, and their realizations $\mathcal{D}:=\{(\bm{x}_{1},y_{1}),\ldots,(\bm{x}_{n},y_{n})\}$ are observed. From these observations, a $p$-dimensional estimate $\widehat{\bm{\beta}}$ is constructed for prediction, with $p$ being larger than the sample size $n$. In the asymptotic setting, where $n$ diverges, $p$ possesses a larger order of divergence, i.e., $p=\omega(n)$.
The primary goal of this study is to examine the generalization performance of a predictor trained on the observations $\mathcal{D}$ in the overparameterized domain. We adopt as the predictor the mean $g(\bm{x}^{\top}\bm{\beta})$, with the estimated parameter $\widehat{\bm{\beta}}$ plugged in. Then, the notion of excess risk is introduced to quantify the generalization performance, measuring the difference between the oracle predictor $g(\bm{X}^{\top}\bm{\beta}^{*})$ and the predictor $g(\bm{X}^{\top}\widehat{\bm{\beta}})$:
$\|g(\bm{X}^{\top}\widehat{\bm{\beta}})-g(\bm{X}^{\top}\bm{\beta}^{*})\|_{P_{\bm{X}},2}.$
where $\|\cdot\|_{P_{\bm{X}},2}$ denotes the $L_{2}(P_{\bm{X}})$ norm with respect to the distribution $P_{\bm{X}}$ of the covariates. As the later analysis shows, the convergence of the excess risk is related to the distributional convergence of the estimated parameter $\widehat{\bm{\beta}}$.
2.2. Method
The predictor construction follows a systematic approach: first, the data is divided into two sets; then, a prior distribution of $\bm{\beta}$ is developed based on the empirical spectral information of $\Sigma$ from one set; next, the likelihood is computed in the other set; finally, the posterior distribution of $\bm{\beta}$ is computed based on the prior and likelihood.
To begin, we divide the dataset $\mathcal{D}$ into two equal parts, $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$, each containing $n/2$ observations. While the split ratio is arbitrary, it is crucial that it does not depend on the sample size $n$. This data splitting allows us to use $\mathcal{D}_{1}$ for constructing the prior distribution and $\mathcal{D}_{2}$ for evaluating the likelihood function, ensuring that the data used in these two steps are independent.
Next, we consider the empirical covariance matrix of the covariates $\bm{X}$, defined as:
$\widehat{\Sigma}:=\frac{2}{n}\sum_{i=1}^{n/2}(\bm{x}_{i}-\bar{\bm{x}})(\bm{x}_{i}-\bar{\bm{x}})^{\top},$ (3)
where $\bar{\bm{x}}\in\mathbb{R}^{p}$ is the sample mean of the covariates from $\mathcal{D}_{1}$. The factor $2/n$ ensures that $\widehat{\Sigma}$ is an unbiased estimator of the population covariance matrix. We then perform a spectral decomposition of $\widehat{\Sigma}$:
$\widehat{\Sigma}=\sum_{j=1}^{p}\widehat{\lambda}_{j}\widehat{\bm{v}}_{j}\widehat{\bm{v}}_{j}^{\top},$ (4)
where $\{\widehat{\lambda}_{j}\}_{j=1}^{p}$ and $\{\widehat{\bm{v}}_{j}\}_{j=1}^{p}$ are the eigenvalues and eigenvectors of $\widehat{\Sigma}$, respectively, with $\widehat{\lambda}_{1}\geq\widehat{\lambda}_{2}\geq\cdots\widehat{\lambda}_{p}\geq 0$. Using the empirical spectra, we define a low-rank approximation of $\widehat{\Sigma}$ as:
$\widehat{\Sigma}_{1:k}:=\sum_{j=1}^{k}\widehat{\lambda}_{j}\widehat{\bm{v}}_{j}\widehat{\bm{v}}_{j}^{\top}.$ (5)
This approximation captures the most significant $k$ eigenvalues and eigenvectors of $\widehat{\Sigma}$, effectively reducing the dimensionality of the covariate space while retaining the most informative directions of variation. The choice of $k$ is a tuning parameter; however, to manage computational cost, we fix it in an interval $[L_{\kappa}:U_{\kappa}]$ such that it satisfies Assumption 2.
We then employ the following prior distribution for $\bm{\beta}$.
Definition 1 (Prior Distribution with Effective Spectra).
$\pi_{\mathfrak{b}}(\bm{\beta}\mid\widehat{\Sigma},k)=\frac{\exp\big{(}-\bm{\beta}^{\top}\widehat{\Sigma}_{1:k}^{\dagger}\bm{\beta}\big{)}\textbf{1}\{\|\bm{\beta}\|_{2}\leq R\}}{\int_{\mathbb{R}^{p}}\exp\big{(}-\bar{\bm{\beta}}^{\top}\widehat{\Sigma}_{1:k}^{\dagger}\bar{\bm{\beta}}\big{)}\textbf{1}\{\|\bar{\bm{\beta}}\|_{2}\leq R\}\mathrm{d}\bar{\bm{\beta}}},$ (6)
where $R>0$ is a fixed constant that determines the radius of the truncation sphere.
Although this prior distribution resembles a $p$-variate Gaussian distribution with mean $\bm{0}$ and covariance $\widehat{\Sigma}_{1:k}$, there are some important distinctions. First, the distribution is truncated, with the indicator function restricting the domain to a centered sphere of radius $R$. This truncation is necessary because the low-rank approximation $\widehat{\Sigma}_{1:k}$ is singular, and defining the density function without support restrictions would cause the integrals over $\mathbb{R}^{p}$ to diverge. Second, unlike [42, 23, 54], we impose a non-sparse prior distribution, even though some parameters may be redundant. This choice is motivated by the desire to retain all potentially relevant covariates in the data and to avoid the challenges associated with sparse priors, such as the need to specify sparsity-inducing hyperparameters.
Next, we consider the posterior distribution, which combines the prior distribution and the likelihood function using Bayes’ Theorem.
Definition 2 (Posterior Distribution).
The probability density function of the posterior distribution is given by:
$\pi(\bm{\beta}\mid\mathcal{D},k)=\frac{\pi_{\mathfrak{b}}(\bm{\beta}\mid\widehat{\Sigma},k){\prod_{(y_{i},\bm{x}_{i})\in\mathcal{D}_{2}}L(y_{i};g(\bm{x}_{i}^{\top}\bm{\beta}))}}{\int_{\mathbb{R}^{p}}\pi_{\mathfrak{b}}(\bm{\beta}\mid\widehat{\Sigma},k){\prod_{(y_{i},\bm{x}_{i})\in\mathcal{D}_{2}}L(y_{i};g(\bm{x}_{i}^{\top}\bm{\beta}))}\mathrm{d}\bm{\beta}},$ (7)
For any measurable set $B\subset\mathbb{R}^{p}$, the posterior probability of $\bm{\beta}\in B$ is given by:
$\Pi\left(\{\bm{\beta}:\bm{\beta}\in B\}\mid\mathcal{D},k\right)=\int_{B}\pi(\bm{\beta}\mid\mathcal{D},k)\mathrm{d}\bm{\beta}.$ (8)
The Bayesian framework allows for the quantification of events of interest, such as the probability of predictions being close to the true values, through the probability measures defined in this manner. Moreover, as will be demonstrated later, the distributional information provided by the posterior is valuable in practice.
Note that we discard the information about the response variable $y_{i}$ in $\mathcal{D}_{1}$ since this subset was only used to compute the sample covariance matrix of the covariates. If dropping this information seems inefficient, one can consider switching the roles of $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ to construct a new estimator (posterior distribution), and averaging the two estimators in a manner similar to the sample splitting scheme [7]. The theoretical results presented later are preserved for this averaging.
3. Main Results
In this section, we present the main theoretical results. We begin by introducing the key assumptions that underlie our analysis (§3.1). These assumptions concern the spectral properties of the covariance matrix and the concentration of the empirical covariance matrix around its population counterpart. We then proceed to establish the posterior contraction for the GLM (§3.2). Our theoretical results offer characterizations of the posterior contraction, elucidating the intricate interplay between the dimensionality, sample size, and the eigenvalue decay of the covariance matrix.
3.1. Assumptions
We begin by focusing on the boundedness of the operator norm of $\Sigma$ and the concentration of the empirical covariance matrix $\widehat{\Sigma}$ around $\Sigma$.
Assumption 1.
The operator norm $\|\Sigma\|_{\mathrm{op}}$ is bounded, and there exist sequences $\{\rho_{n}\}_{n\in\mathbb{N}}$ and $\{\ell_{n}\}_{n\in\mathbb{N}}$ of non-negative reals with $\rho_{n},\ell_{n}=o(1)$ as $n\to\infty$, such that for any sufficiently large $n\in\mathbb{N}$,
$\mathbb{P}\left(\frac{\|\widehat{\Sigma}-\Sigma\|_{\mathrm{op}}}{\|\Sigma\|_{\mathrm{op}}}\leq\rho_{n}\right)\geq 1-\ell_{n}.$ (9)
This assumption ensures that the empirical covariance matrix concentrates well around the population covariance matrix. The sequence $\rho_{n}$ quantifies the rate of this concentration, while $\ell_{n}$ controls the probability of the concentration event. It is worth noting that this assumption is mild and includes not only the sub-Gaussian distributions commonly imposed in high-dimensional statistics but also log-concave distributions, which encompass most unimodal parametric distributions [56, 57] and allow for heavy-tailed distributions such as the Laplace distribution. In contrast, other related literature has imposed more restrictive conditions such as sub-Gaussianity [5] or strong log-concavity [13]. In a high-dimensional setting, where many covariates are incorporated, the mildness of the distributional assumption is particularly important, as it allows for a wider range of applicable scenarios.
Our next assumption concerns the behavior of the eigenvalues of $\Sigma$.
Assumption 2.
Given a covariance matrix $\Sigma$ satisfying (9), with eigenvalues $\{\lambda_{k}\}_{k=1}^{p}$, the sequence $\{\rho_{n}\}_{n\in\mathbb{N}}$ from Assumption 1, and a sequence of bounds $(L_{\kappa},U_{\kappa})$, there exists a sequence $\{\varepsilon_{n}\}_{n\in\mathbb{N}}$ of positive reals such that the following hold as $n\to\infty$:
$\mathrm{(i)}$ $\quad\varepsilon_{n}\to 0,\quad n^{-1/2}\lesssim\rho_{n}\lesssim\varepsilon_{n}^{2},\qquad\mathrm{(ii)}\quad(\rho_{n}+\lambda_{k})\lambda_{k}\lessapprox\frac{\varepsilon_{n}^{2}}{k},\quad\forall k\in[L{\kappa}:U_{\kappa}],$ (10)
$\mathrm{(iii)}$ $\quad\sum_{k\in[L_{\kappa}+1:p]}\lambda_{k}\lesssim\varepsilon_{n}^{4},\qquad\mathrm{(iv)}\quad U_{\kappa}\lessapprox n\varepsilon^{2}.$ (11)
This assumption imposes certain regularity conditions on the eigenvalues of $\Sigma$. Intuitively, conditions (i) and (ii) control the rate at which the bulk of the eigenvalues (those indexed from $L_{\kappa}$ to $U_{\kappa}$) are allowed to decay moderately. Condition (iii) requires that the eigenvalues beyond the bulk (those indexed larger than $U_{\kappa}$) are sufficiently small, while condition (iv) constrains the size of $U_{\kappa}$ relative to the sample size $n$ and the rate $\varepsilon_{n}$. These conditions are crucial for deriving the posterior contraction rates in Theorem 1.
Finally, we assume a certain local Lipschitz condition of the link function $g$ in terms of the Kullback-Leibler (KL) divergence and the variational form between the corresponding probability distributions.
Assumption 3.
For any pair $\bm{\beta},\bm{\beta}^{\prime}\in\mathbb{R}^{p}$, the GLM with link function $g$ satisfies
$K(P,P^{\prime}),~{}V(P,P^{\prime})\lesssim\|\bm{\beta}-\bm{\beta}^{\prime}\|_{\Sigma}^{2},$
where $P$ and $P^{\prime}$ are the probability distributions corresponding to the densities $L(y;g(\bm{x}^{\top}\bm{\beta}))$ and $L(y;g(\bm{x}^{\top}\bm{\beta}^{\prime}))$, respectively.
This assumption limits the amount of variation in a distribution when different parameters are used, meaning that it controls the complexity of the model. This assumption is satisfied by several common GLMs, as illustrated in the following examples.
Example 1.
  1. (1)
    For normal linear regression with known variance $\sigma^{2}$, we have
    $2K(P,P^{\prime})=V(P,P^{\prime})\leq\|g(\bm{X}^{\top}\bm{\beta})-g(\bm{X}^{\top}\bm{\beta}^{\prime})\|_{P_{\bm{X}},2}^{2}\lesssim\|\bm{\beta}-\bm{\beta}^{\prime}\|_{\Sigma}^{2},$
    where the last inequality holds if the link function $g$ is Lipschitz continuous.
  2. (2)
    For binary regression with the logistic link, we have
    $K(P,P^{\prime}),V(P,P^{\prime})$ $\leq\|\bm{\beta}-\bm{\beta}^{\prime}\|_{\Sigma}^{2},$
    which are both bounded by $\|\bm{\beta}-\bm{\beta}^{\prime}\|_{\Sigma}$. In contrast, for the probit link, Assumption 3 does not hold in general.
  3. (3)
    For Poisson regression with a bounded and positive link function $g$, we have
    $K(P,P^{\prime}),V(P,P^{\prime})$ $\lesssim\|\bm{\beta}-\bm{\beta}^{\prime}\|_{\Sigma}^{2}.$
    Although the exponential function, $g(r)=e^{r}$, is the canonical link function for Poisson regression, it is unbounded, and Assumption 3 does not hold.
These examples demonstrate that Assumption 3 is a reasonable condition that holds for several widely used GLMs.
3.2. Posterior Contraction
We now present the main result, which establishes the posterior contraction for the GLM under the assumptions introduced in the previous subsection.
Theorem 1.
Consider the GLM (2) with covariate distribution satisfying Assumptions 1-3, and a prior distribution on $\bm{\beta}$ is (6). Then, for a sufficiently large constant $M$, the posterior distribution contracts around the true parameter $\bm{\beta}^{*}$ as
[…]
Furthermore, if the eigenvalues of $\Sigma$ satisfy the additional condition $(nk)^{-1}\gtrsim(\rho_{n}+\lambda_{k})\lambda_{k}$ for all $k\in[L_{\kappa}:U_{\kappa}]$, then the contraction rate can be explicitly characterized as
$\varepsilon_{n}=O\left(U_{\kappa}\lambda_{L_{\kappa}}^{1/2}(\lambda_{L_{\kappa}}+\rho_{n})^{1/2}\right).$ (13)
This theorem shows that under suitable conditions on the covariance matrix $\Sigma$ and the prior distribution, the posterior distribution of the predictor concentrates around the oracle predictor at a rate $\varepsilon_{n}$ with respect to the $L_{2}(P_{\bm{X}})$ distance. The rate $\varepsilon_{n}$ depends on the interplay between the sample size $n$, the concentration rate $\rho_{n}$ of the empirical covariance matrix, and the eigenvalues of $\Sigma$ in the bulk region.
3.2.1. Consistency as a distribution
While Theorem 1 derives posterior consistency of the predictor with respect to the $L_{2}$ distance, it also holds for other commonly used distances between probability distributions, such as the total variation distance, the Hellinger distance, and the KL divergence in several cases, including a list of distributions in Example 1. The contraction in these divergences is particularly useful for assessing the quality of the posterior distribution in terms of its ability to approximate the true data-generating distribution. A detailed discussion of these results is provided in the appendix.
3.2.2. Consistency of Point Estimators
The posterior contraction result in Theorem 1 not only provides a theoretical justification for using the posterior distribution as a reliable approximation of the true data-generating distribution but also has important implications for point estimators in the usual sense of consistency of estimators. Two commonly used point estimators in Bayesian inference are the maximum a posteriori (MAP) estimator and the posterior mean.
The MAP estimator, denoted as $\widehat{\bm{\beta}}_{\text{MAP}}$, is defined as the mode of the posterior distribution:
$\widehat{\bm{\beta}}_{\text{MAP}}=\operatornamewithlimits{argmax}_{\bm{\beta}\in\mathbb{R}^{p}}\Pi(\bm{\beta}\mid\mathcal{D}).$
Under the same assumptions as in Theorem 1, it can be shown that the MAP estimator is consistent with respect to the $L_{2}(P_{\bm{X}})$ norm:
$\|g(\bm{X}^{\top}\widehat{\bm{\beta}}_{\text{MAP}})-g(\bm{X}^{\top}\bm{\beta}^{*})\|_{P_{\bm{X}},2}\overset{P^{*}}{\longrightarrow}0,\quad\mbox{as}~{}n\to\infty.$
Another widely used point estimator is the posterior mean, denoted as $\widehat{\bm{\beta}}_{\text{PM}}$, which is defined as the expectation of the parameter with respect to the posterior distribution:
[…]
Under the assumptions of Theorem 1 and the additional condition that posterior consistency with respect to the total variation distance or the Hellinger distance holds (as discussed in §3.2.1), the posterior mean is also consistent:
$\|g(\bm{X}^{\top}\widehat{\bm{\beta}}_{\text{PM}})-g(\bm{X}^{\top}\bm{\beta}^{*})\|_{P_{\bm{X}},2}\overset{P^{*}}{\longrightarrow}0,\quad\mbox{as}~{}n\to\infty.$ (15)
As the posterior mean considers the weights of the entire distribution (as it is defined by integral), its consistency requires the additional condition that the posterior distribution is not ill-behaved. We defer the further discussion to the appendix.
4. Extension to Two-Parameter Gaussian Case
In this section, we extend our analysis to the setting where the GLM has a Gaussian distribution for the measurement error. Specifically, we consider the following model, referred to as the single neuron model [47, 60]:
$Y=g(\bm{X}^{\top}\bm{\beta})+\varepsilon;\quad\varepsilon\sim N(0,\sigma^{2}),$ (16)
where $\sigma>0$ is the standard deviation of the error term. This model is more flexible than the standard linear regression model and can accommodate various link functions $g$. For instance, if $g$ is a censored link function such as the ReLU, tanh, or softplus, then (16) represents a censored regression model.
Here, we specify the prior distributions for the parameters $\bm{\beta}$ and $\sigma^{2}$. For $\bm{\beta}$, we continue to use the prior distribution (6). Regarding the variance parameter, we employ as the prior distribution the inverse Gaussian distribution $\Pi_{\varsigma^{2}}(\cdot)$ with the density of the form:
$\pi_{\varsigma^{2}}(\sigma^{2})=\sqrt{\frac{\xi}{2\pi\sigma^{6}}}\exp\left(\frac{-\xi(\sigma^{2}-\eta)}{2\eta^{2}\sigma^{2}}\right),$ (17)
where $\eta>0$ and $\xi>0$ denote the mean and shape parameters, respectively. We allow for arbitrary assignment of hyperparameters. An important property of the inverse Gaussian prior is that the tails on both sides decrease rapidly, which is critical for posterior contraction. In contrast, the inverse gamma distribution, the conjugate prior to the Gaussian likelihood, has a light lower tail and a heavy upper tail. From a Bayesian perspective, the inverse Gaussian prior reflects the analyst’s belief that the parameter will not have extremely large or small values. This prior distribution for the variance parameter has also been employed in [51, 41, 59].
We now present an analog of Theorem 1 for the two-parameter Gaussian model (16).
Theorem 2.
Consider the regression model (16) with covariate distribution satisfying Assumptions 1-3, and a prior distribution on $(\bm{\beta},\sigma^{2})$ satisfying (6) and (17). Then, for a sufficiently large constant $M$, the following convergences hold as $n\to\infty$:
$\Pi\Big{(}\{\bm{\beta}:\|g(\bm{X}^{\top}\bm{\beta})-g(\bm{X}^{\top}\bm{\beta}^{*})\|_{P_{\bm{X}},2}>M\varepsilon_{n}\}\mid\,\mathcal{D}\Big{)}$ $\overset{P^{*}}{\longrightarrow}0,$ (18)
$\Pi\Big{(}\{\sigma^{2}:|\sigma^{2}-(\sigma^{*})^{2}|>M\varepsilon_{n}\}\,\mid\,\mathcal{D}\Big{)}$ $\overset{P^{*}}{\longrightarrow}0,$ (19)
where $(\bm{\beta}^{*},\sigma^{*})$ denotes the true parameter pair. Furthermore, if the eigenvalues of $\Sigma$ satisfy the additional condition $(nk)^{-1}\gtrsim(\rho_{n}+\lambda_{k})\lambda_{k}$ for all $k\in[L_{\kappa}:U_{\kappa}]$, then the contraction rate can be explicitly characterized as
$\varepsilon_{n}=O\left(U_{\kappa}\lambda_{L_{\kappa}}^{1/2}(\lambda_{L_{\kappa}}+\rho_{n})^{1/2}\right).$ (20)
This theorem provides the posterior contraction for both the predictor $g(\bm{X}^{\top}\bm{\beta})$ and the error variance $\sigma^{2}$ in the two-parameter Gaussian model. The contraction rate for the predictor is with respect to the $L_{2}(P_{\bm{X}})$ distance, while the rate for $\sigma^{2}$ is with respect to the absolute difference.
The extension to the two-parameter Gaussian model is particularly relevant for applications where the error variance is unknown and needs to be estimated from the data. By considering the joint posterior contraction rates for $(\bm{\beta},\sigma^{2})$, Theorem 2 presents a theoretical justification for Bayesian inference in such models and offers guidance on the choice of prior distributions that can adapt to the unknown error variance.
5. Simulation
This section confirms by simulation the statement of the theorem presented in §3 and §4.
5.1. Logistic Regression
Here, we consider a binary classification scenario where the response variable $Y_{i}$ takes values in $\{0,1\}$. The response variables are generated according to the following model:
$Y_{i}=\mathrm{Ber}(g_{ls}(\bm{X}_{i}^{\top}\bm{\beta}));\quad i=1,2,...,n,$
where $g_{ls}(r):=1/(1+e^{-r})$ is the logistic function. We set $\bm{\beta}\sim N(\bm{0},I_{p})$ and $p=n^{4/3}$, ensuring an overparameterized setting. Regarding the covariate, we consider two settings: (A) Gaussian covariate, that is, $\bm{X}_{i}\sim N(\bm{0},\Sigma)$; (B) Laplacian convariate, that is, $\bm{X}_{i}\sim L(\bm{0},\Sigma)$, where the eigenvalues of $\Sigma$ follow $\lambda_{j}(\Sigma)=10\exp(-j/8)+n\exp\{-\sqrt{n}\}/p$.
We investigate the behavior of the proposed method as the sample size increases by varying the training data size $n$ from $50$ to $500$. For each sample size, we generate a test dataset of $1000$ instances to evaluate the performance of the trained classifier. To evaluate the performance of the Bayesian predictions, we employ four widely used metrics: 0-1 loss, which quantifies the misclassification rate; Area Under the ROC Curve (AUC), which assesses the discriminative ability of the classifier; Unconfident Misclassifications rate (UM), which measures the proportion of unconfident predictions among misclassified instances; and Confident Correct predictions rate (CC), which calculates the proportion of accurate predictions among confident ones. We refer to a confident prediction as one in which the $95\%$ interval does not cross the threshold $(0.5)$, while an unconfident one does. The latter two metrics are helpful when building systems that defer action when there is no confidence in the prediction.
Figure 1. Change of 0-1 loss, AUC (area under the ROC curve), UM (proportion of unconfident predictions among misclassifications), and CC (proportion of accurate predictions among confident ones) as the sample size grows in the logistic regression with the Gaussian covariate case (left) and the Laplace covariate case (right).
Figure 1 illustrates the results of the experiment. As the sample size increases, the 0-1 loss exhibits a consistent decrease in both settings, indicating an improvement in classification accuracy. This trend persists even in the overparameterized setting, aligning with the theoretical findings. The AUC metric also shows a steady increase with the sample size, suggesting an enhanced ability to discriminate between the two classes. The UM and CC metrics provide insights into the uncertainty estimation of the proposed method. UM consistently presents a high value. This indicates that the method is able to warn us of a lack of confidence when misclassifying. The CC also presents a similarly high value, which means that confident predictions are trustworthy. These observations suggest that the proposed method effectively captures the uncertainty associated with predictions and leverages it to make more informed decisions.
5.2. Nonlinear regression with Gaussian noise
In this subsection, we consider a scenario where the true response takes values in the positive real line $\mathbb{R}_{+}$. Specifically, the response variables are generated as follows:
$Y_{i}=N(g_{sp}(\bm{X}_{i}^{\top}\bm{\beta}),1);\quad i=1,2,...,n,$
where $\bm{\beta}\sim N(\bm{0},I_{p})$, $p=n^{4/3}$, and $g_{sp}(r):=\log(1+e^{r})$ is the softplus function. The distributions of the covariate vectors are set to follow the same patterns as in the previous experiment.
We examine how the performance of Bayesian predictions changes as the data size increases by varying the training data size $n$ from $50$ to $500$. For a newly generated test dataset $\{(\bm{x}_{i},g_{sp}(\bm{x}_{i}^{\top}\bm{\beta}))\}_{i=1}^{1000}$, we assess the accuracy of Bayesian predictions using three widely adopted indicators: Root Mean Squared Error (RMSE), measuring the average deviation between the predicted values and the true values; Coverage Probability (CP), evaluating the reliability of the $95\%$ prediction intervals by calculating the proportion of true values falling within the intervals; and Average Length (AL), quantifying the width of the $95\%$ prediction intervals, providing insight into the precision of the predictions.
Figure 2. Change of RMSE (root mean squared error), CP (coverage probability of $95$% prediction interval), and AL (average length of $95$% prediction interval) as the sample size grows in nonlinear Gaussian regression with the Gaussian covariate case (left) and the Laplace covariate case (right).
The results are shown in Figure 2. First, the RMSE decreases monotonically as the sample size increases. Even when $p$ becomes large and the setting becomes clearly overparameterized, the fact that the accuracy continues to improve supports the theoretical results experimentally. Next, we focus on CP and AL. These are indicators related to prediction intervals, with CP representing the reliability of the prediction interval and AL representing its width. From the figure, CP is not good when $n=100$, but after that, it stabilizes at a high value. AL is also slightly increasing, but not too much, implying that the interval is not trivial. (Regardless of $n$, the mean of $\{g_{sp}(\bm{x}_{i}^{\top}\bm{\beta})\}_{i=1}^{1000}$ in the test data is around $3.5$ and the variance is around $5$.) These results suggest that the proposed method is able to properly quantify the uncertainty in constructing prediction intervals.
6. Data Analysis
6.1. Dataset
We apply the proposed method to the ARCENE dataset, which was originally curated for the NIPS 2003 feature selection challenge for the binary classification problem. The dataset consists of mass spectrometric measurements from cancer patients and healthy individuals, with the goal of distinguishing between the two groups based on these measurements. The ARCENE dataset, accessible through the UC Irvine Machine Learning Repository [16], was constructed by merging three separate datasets: National Cancer Institute (NCI) ovarian cancer data; NCI prostate cancer data; and Eastern Virginia Medical School (EVMS) prostate cancer data. In total, the public dataset contains $200$ samples, comprising $88$ cancer samples and $112$ normal samples. These data are categorized as cancer samples and control samples, with $10000$ continuous covariates, with 7000 dimensions corresponding to real mass-spectrometry peaks and $3000$ dimensions representing random noise variables (probes). The primary challenges posed by the dataset lie in its high dimensionality and the presence of overlapping classes. The goal here is to discriminate between cancer and normal samples.
6.2. Experimental Setting
We compared the performance of five distinct methods.
  • Logistic Regression with the effective spectral (ES) prior distribution: The proposed Bayesian methodology, as described in Section 2.2, was applied to the logistic regression model to make predictions on unseen data.
  • Logistic Regression with the Horseshoe (HS) prior: To handle the high-dimensionality of covariates, we implemented logistic regression with the Horseshoe prior [12]. The implementation was carried out using the ”horseshoenlm” package in R [38].
  • Logistic Regression with the Normal (NM) prior: We investigated the performance of logistic regression with a standard normal prior distribution, which is the most commonly used prior and serves as a ridge-like regularization [15]. This prior allows for flexibility in modeling the coefficient distribution and can be easily implemented using the Pyro probabilistic programming language [3].
  • Logistic Regression with the Laplace (LP) prior: We also explored the performance of logistic regression with a standard Laplace prior distribution, which promotes sparsity in the coefficient estimates as a Bayesian Lasso [43]. This prior has been shown to be effective in high-dimensional settings and can be efficiently implemented using Pyro, leveraging its support for non-Gaussian prior distributions [3].
  • LightGBM: As a baseline method for table data prediction tasks, we implemented LightGBM [27], an efficient and scalable gradient boosting framework. We adopted the ”lightgbm” package in Python for the implementation and trained the model using Bayesian optimization [48, 14] via the ”Optuna” package [1].
For all methods, we conducted performance evaluation using 5-fold cross-validation. In each fold, the model was trained on the training data, and predictions were made on the test data. As in the previous section, the evaluation metrics used to assess predictive performance were Accuracy, Area Under the ROC Curve (AUC), UM, and CC.
6.3. Results
The results presented in Table 1 provide valuable insights into the performance of the proposed Bayesian logistic regression approach with different prior distributions, as well as its comparison to the LightGBM baseline. Overall, Logistic Regression with the ES prior distribution achieves the highest accuracy and AUC among all the methods, demonstrating its effectiveness in capturing the underlying structure of the high-dimensional ARCENE dataset. The superior performance of the ES prior can be attributed to its ability to adapt to the intrinsic dimensionality of the data by incorporating the empirical spectral information of the covariance matrix [59].
Let us take a closer look at the comparison. The horseshoe prior, known for its flexibility and sparsity-inducing properties [12], also performs well, with the second-highest accuracy and AUC. This suggests that the regression on ARCENE dataset may exhibit some degree of sparsity, which the HS prior is able to capture effectively. However, the ES prior outperforms the HS prior, indicating that the adaptive nature of the ES prior is more beneficial in this context. Also, the fact that the $3000$ covariates are essentially irrelevant to the regression and have zero coefficients but are not highly sparse, may be the reason why the proposed method prevails. The LP and NM priors, while still providing competitive results, do not perform as well as the ES and HS priors. Nevertheless, their performance is still superior to the LightGBM baseline, highlighting the advantages of the Bayesian approach in handling excess covariates.
A notable aspect of the results is the uncertainty estimation of the Bayesian methods. The ES prior achieves the highest UM and CC values, indicating that it is highly effective in identifying uncertainty and providing reliable confidence estimates. This is particularly valuable in practical applications, where the ability to defer decision-making when the model is uncertain can prevent the risk of misclassifications. The HS prior also demonstrates good uncertainty quantification, with the second-highest UM value. However, the LP and NM priors exhibit very low UM values, suggesting that they may be overconfident in their predictions. Note that the LightGBM method does not provide uncertainty estimation, as it is not a probabilistic model.
Model Accuracy AUC UM CC
Logistic with ES 0.895 0.950 0.860 0.963
Logistic with HS 0.845 0.900 0.632 0.908
Logistic with LP 0.805 0.886 0.009 0.936
Logistic with NM 0.815 0.883 0.008 0.936
LightGBM 0.775 0.858
Table 1. Data Analysis: Data Analysis: Comparison of classification performance between our method, LightGBM. The metrics are Accuracy, Area Under the Curve (AUC), the proportion of uncertain predictions among misclassifications (UM), and the proportion of correct predictions among certain classifications (CC).
7. Discussion
This study introduces a novel Bayesian approach for prediction in overparameterized generalized linear models, highlighting its theoretical foundations and practical applications. The primary theoretical contribution lies in establishing posterior contraction under appropriate conditions, validating the use of the posterior distribution as a reliable approximation of the true data-generating distribution. Empirical evidence from simulations and real-world datasets corroborates the effectiveness of the proposed method in delivering accurate predictions and reliable uncertainty estimates.
In the realm of Bayesian high-dimensional inference, the proposed approach distinguishes itself by directly addressing non-sparsity while offering theoretical guarantees for prediction with nonlinearity. This unique perspective paves the way for new avenues of Bayesian inference in overparameterized models.
This study marks a significant advance in overparameterization theory by achieving consistency in nonlinear regression problems. The robustness of posterior contraction theory to model complexity (nonlinearity) plays a pivotal role in attaining consistency because it solely necessitates control of the covering number, the decay of the prior distribution in its tail region, and the concentration of the prior distribution around the true value of the parameter.
Finally, we conclude this paper by discussing its limitations and identifying potential areas for improvement. The theoretical analysis hinges on specific assumptions regarding the covariance matrix, which may not always hold in real-world scenarios. Future research could focus on relaxing these assumptions and extending the results to more general settings as deep neural networks generalize to a wide range of data. Moreover, the computational complexity associated with sampling from the posterior distribution can pose challenges for large-scale datasets. Also, further improvements in uncertainty estimation could also be combined with the more flexible approximation such as normalizing flow [44, 45] and Stein Variational Gradient Descent [35].
Acknowledgements
I would like to express my sincere gratitude to my advisor, Masaaki Imaizumi, for his patient and rigorous supervision. This work was supported by JSPS KAKENHI (22J21090) and JST ACT-X (JPMJAX23CS).
Appendix A Proof of the Main Results
In this section, we provide proof of the main results presented in the previous sections. The key is to establish the posterior contraction rate for the squared $L_{2}(P_{\bm{X}})$ norm of the difference between the true and estimated mean functions, i.e., $\mathbb{E}[\{g(\bm{X}^{\top}\widehat{\bm{\beta}})-g(\bm{X}^{\top}\bm{\beta}^{*})\}^{2}]$. Since the link function $g$ is assumed to be Lipschitz continuous, we have the following upper bound (up to a constant factor):
$\|g(\bm{X}^{\top}\widehat{\bm{\beta}})-g(\bm{X}^{\top}\bm{\beta}^{*})\|_{P_{\bm{X}},2}\lesssim\|\widehat{\bm{\beta}}-\bm{\beta}^{*}\|_{\Sigma},$ (21)
where $\|\cdot\|_{\Sigma}$ denotes the norm induced by the covariance matrix $\Sigma$. This bound relates the contraction of the posterior distribution in the $L_{2}(P_{\bm{X}})$ norm to the contraction in the $\|\cdot\|_{\Sigma}$ norm, which is similar to the predictive risk in linear regression models [5, 39, 11]. Therefore, it suffices to show that the posterior distribution of $\|\widehat{\bm{\beta}}-\bm{\beta}^{*}\|_{\Sigma}$ concentrates around zero at a certain rate.
To this end, we invoke a general result on posterior contraction rates from [17]. Let $\mathcal{P}=\mathbb{R}^{p}$ be the parameter space equipped with the distance $d_{\mathcal{P}}$ induced by the $\|\cdot\|_{\Sigma}$ norm, and define a sequence of restricted parameter spaces
$\mathcal{P}_{n}:=\{\bm{h}\in\mathbb{R}^{p}:\|\bm{h}\|_{2}\leq H_{n}\},$ (22)
where $H_{n}>0$ is a radius satisfying $H_{n}\to\infty$ as $n\to\infty$. Recall from Assumption 1 that the empirical covariance matrix $\widehat{\Sigma}$ concentrates around $\Sigma$ with probability at least $1-\ell_{n}$, which controls the randomness of the posterior distribution.
Theorem 2.1 of [17] states that the posterior contraction rate in Theorem 1 holds if there exists a sequence $H_{n}\to\infty$ such that the following three conditions are satisfied for any sufficiently large $n$ and some constant $C>0$:
$\log\mathcal{N}(\varepsilon_{n},\mathcal{P}_{n},d_{\mathcal{P}})\leq n\varepsilon_{n}^{2},$ (23)
$\Pi_{\mathfrak{b}}(\mathcal{P}\setminus\mathcal{P}_{n}\mid\widehat{\Sigma})\leq\exp(-(C+4)n\varepsilon_{n}^{2}),$ (24)
$\Pi_{\mathfrak{b}}\left(\left\{\bm{\beta}:K(P^{*},P_{\bm{\beta}})\leq\varepsilon_{n}^{2},V(P^{*},P_{\bm{\beta}})\leq\varepsilon_{n}^{2}\right\}\mid\widehat{\Sigma}\right)\geq\exp(-Cn\varepsilon_{n}^{2}).$ (25)
The first two conditions (23) and (24) are determined by the tuple $(\mathcal{P},\mathcal{P}_{n},d_{\mathcal{P}},\Pi_{\mathfrak{b}})$. Since these conditions have been verified in [59] for the same choice of the parameter spaces and prior distribution, we omit their proofs here.
To establish the remaining condition (25), we observe that by Assumption 3,
$\Pi_{\mathfrak{b}}\left(\left\{\bm{\beta}:K(P^{*},P_{\bm{\beta}})\leq\varepsilon_{n}^{2},V(P^{*},P_{\bm{\beta}})\leq\varepsilon_{n}^{2}\right\}\mid\widehat{\Sigma}\right)\geq\Pi_{\mathfrak{b}}\left(\left\{\bm{\beta}:\|\bm{\beta}-\bm{\beta}^{*}\|_{\Sigma}^{2}\leq\varepsilon_{n}^{2}\right\}\mid\widehat{\Sigma}\right).$
Therefore, it suffices to show that
$\Pi_{\mathfrak{b}}\left(\left\{\bm{\beta}:\|\bm{\beta}-\bm{\beta}^{*}\|_{\Sigma}^{2}\leq\varepsilon_{n}^{2}\right\}\mid\widehat{\Sigma}\right)\geq\exp(-Cn\varepsilon_{n}^{2}).$ (26)
The proof of (26) follows from a similar argument as in [59], which we outline below for completeness.
The extension to the case where $\sigma^{2}$ is unknown is straightforward. If we take $[1/n,n]$ as $\mathcal{P}_{n}$ when we have an inverse Gaussian distribution as a prior distribution for $\sigma^{2}$, the posterior contraction holds because $\mathcal{P}_{n}$ dominates $\mathcal{P}$ as $n$ increases and the remaining part $\mathcal{P}\setminus\mathcal{P}_{n}$ has only the negligible mass of the prior.
Combining the above arguments, we conclude that the conditions of Theorem 2.1 in [17] are satisfied, and hence the posterior contraction in Theorem 1 holds. The explicit characterization of the contraction rate $\varepsilon_{n}$ in terms of the eigenvalues of $\Sigma$ follows from a careful analysis of the interplay between the rate $\rho_{n}$ in Assumption 1, the upper and lower bounds $(L_{\kappa},U_{\kappa})$ in Assumption 2, and the additional assumption on the eigenvalue decay stated in Theorem 1.
Appendix B Consistency as a Distribution and Point Estimator
Following the discussion on posterior distribution contraction, we further explore the convergence of distributions and estimators using various metrics.
To extend the discussion, let us revisit Example 1.
Example 2.
Recall that $K(P,P^{\prime})$ is the KL divergence, $V(P,P^{\prime})$ is the KL variation, and $H(P,P^{\prime})$ is the Hellinger distance between models with different parameters.
  1. (1)
    For normal linear regression with known variance $\sigma^{2}$, we have
    $2K(P,P^{\prime})=V(P,P^{\prime})=H(P,P^{\prime})\leq\|\bm{\beta}-\bm{\beta}^{\prime}\|_{\Sigma}^{2}.$
  2. (2)
    For binary regression with the logistic link, we have
    $K(P,P^{\prime}),V(P,P^{\prime}),H(P,P^{\prime})\lesssim\|\bm{\beta}-\bm{\beta}^{\prime}\|_{\Sigma}^{2}.$
  3. (3)
    For Poisson regression with a link function $g$, we have
    $K(P,P^{\prime})$ $\lesssim C_{K}\|\bm{\beta}-\bm{\beta}^{\prime}\|_{\Sigma}^{2},$
    $V(P,P^{\prime})$ $\lesssim C_{V}\|\bm{\beta}-\bm{\beta}^{\prime}\|_{\Sigma}^{2},$
    $H(P,P^{\prime})$ $\lesssim C_{H}\|\bm{\beta}-\bm{\beta}^{\prime}\|_{\Sigma}^{2},$
    where $C_{K}=C_{\bm{\beta}}\|g^{\prime}/g\|_{\infty}^{2}+C_{\bm{\beta}}\|g^{\prime\prime}/g\|_{\infty}+\|g^{\prime\prime}\|_{\infty}$, $C_{V}=C_{\bm{\beta}}\|g^{\prime}/g\|_{\infty}^{2}$, $C_{H}=\|g^{\prime}/\sqrt{g}\|_{\infty}$ and $C_{\bm{\beta}}=\sup_{\bm{x}}|g(\bm{\beta}^{\top}\bm{x})|$. Note that the factors, $C_{K},C_{V},$ and $C_{H}$, are not necessarily bounded.
As shown in the previous section, the consistency of $\|\bm{\beta}-\bm{\beta}^{\prime}\|_{\Sigma}^{2}$ holds. Consequently, in the above examples, if the distance between distributions can be bounded by constant times the quantity, we obtain the consistency of the posterior distribution to the true data-generating distribution with respect to the KL divergence, KL variation, and Hellinger distance. Furthermore, because the total variation distance and the Hellinger distance induce the same topology, posterior consistency also holds in terms of the total variation distance.
In connection with distributional convergence, we now shift our focus to the convergence of the posterior mean. If the parameter space $\mathcal{P}$ is convex and the distance $d_{\mathcal{P}}$ on $\mathcal{P}$ is bounded and convex, then the posterior mean serves as a consistent estimator ([e.g., Theorem 6.8 in 20]). The assumption of $\mathcal{P}$ being convex is often satisfied, and the boundedness and convexity of the distance hold, for instance, with the total variation distance. Hence, in the above examples, the consistency of the posterior mean is achieved.
Appendix C Computational Method
In this section, we describe the computational method for the proposed approach. We employ stochastic variational inference [19, 21, 4] to approximate the posterior distribution of the model parameters. Since the closed-form expression of the KL divergence is not available, we use the reparameterization trick [28, 46] to enable efficient gradient-based optimization. The reparameterization trick allows us to express the variational distribution in terms of a deterministic function of the variational parameters and a random variable, making the gradients of the variational parameters computable. The optimization part is performed using the Adam algorithm [25].
C.1. Logistic Regression Example
To illustrate the computational procedure, we consider the example of logistic regression. The model components are as follows:
  1. (1)
    Likelihood: $\prod_{(y_{i},\bm{x}_{i})\in\mathcal{D}_{2}}L(y_{i};\bm{x}_{i}^{\top}\bm{\beta})$
  2. (2)
    Prior: $p_{pri}=\phi(\bm{\beta}\mid\bm{0},\widehat{\Sigma}_{1:k})$
  3. (3)
    Posterior: $p_{pos}\propto\phi(\bm{\beta}\mid\bm{0},\widehat{\Sigma}_{1:k})\prod_{(y_{i},\bm{x}_{i})\in\mathcal{D}_{2}}\phi(y_{i};\bm{x}_{i}^{\top}\bm{\beta},\sigma^{2})$
  4. (4)
    Approximator: $q(\bm{\beta}\mid{\widetilde{\bm{\mu}}},\widetilde{\Sigma})$
Indeed, the prior distribution is not just a normal distribution but has a support constraint. However, it can be approximated as a normal distribution if $R$ is taken large enough. Also, a procedure to compute the posterior distribution without support constraints and then project it onto the constrained region is justified by [49].
Our goal is to find the optimal variational parameters ${\widetilde{\bm{\mu}}}_{opt}$ and $\widetilde{\Sigma}_{opt}$ that minimize the KL divergence between the variational approximation $q(\bm{\beta}\mid{\widetilde{\bm{\mu}}},\widetilde{\Sigma})$ and the true posterior $p_{pos}$:
${\widetilde{\bm{\mu}}}_{opt},\widetilde{\Sigma}_{opt}:=\operatornamewithlimits{argmin}_{{\widetilde{\bm{\mu}}},\widetilde{\Sigma}}\mathrm{KL}(q(\bm{\beta}\mid{\widetilde{\bm{\mu}}},\widetilde{\Sigma})\|p_{pos}),$ (27)
where $\mathrm{KL}(p\|q)$ denotes the KL divergence between two probability measures with densities $p$ and $q$. The KL divergence can be expressed as
$\mathrm{KL}(q(\bm{\beta}\mid{\widetilde{\bm{\mu}}},\widetilde{\Sigma})\|p_{pos})=\mathbb{E}_{q}[\log q]-\mathbb{E}_{q}[\log\phi(\bm{\beta}\mid\bm{0},\widehat{\Sigma}_{1:k})]-\sum_{i=1}^{n}\mathbb{E}_{q}[\log L(y_{i};\bm{x}_{i}^{\top}\bm{\beta})]+\mathrm{const}.$ (28)
To compute the gradients of the KL divergence with respect to the variational parameters, we employ the reparameterization trick. Specifically, we denote by $\widetilde{\bm{\beta}}$ a sample from the variational distribution $q$ and parameterize it as $\widetilde{\bm{\beta}}=\widetilde{\bm{\mu}}+\widetilde{\Sigma}_{1/2}\widetilde{\bm{\epsilon}}$, where $\widetilde{\bm{\epsilon}}$ is a sample from a standard Gaussian distribution $\mathcal{N}(\bm{0},\bm{I}_{p})$ and $\widetilde{\Sigma}_{1/2}$ is a decomposition of $\widetilde{\Sigma}$, i.e., $\widetilde{\Sigma}=\widetilde{\Sigma}_{1/2}\widetilde{\Sigma}_{1/2}^{\top}$. With this reparameterization, the KL divergence can be approximated as an unbiased estimate
$\mathrm{KL}(q(\bm{\beta}\mid{\widetilde{\bm{\mu}}},\widetilde{\Sigma})\|p_{pos})\approx\log q(\widetilde{\bm{\beta}}\mid{\widetilde{\bm{\mu}}},\widetilde{\Sigma})-\log\phi(\widetilde{\bm{\beta}}\mid\bm{0},\widehat{\Sigma}_{1:k})-\sum_{i=1}^{n}\log L(y_{i};\bm{x}_{i}^{\top}\widetilde{\bm{\beta}})+\mathrm{const}.$ (29)
We can now compute the gradients of the KL divergence with respect to the variational parameters. The gradient with respect to $\widetilde{\bm{\mu}}$ is given by
$\frac{\partial}{\partial\widetilde{\bm{\mu}}}\mathrm{KL}(q(\widetilde{\bm{\beta}}\mid{\widetilde{\bm{\mu}}},\widetilde{\Sigma})\|p_{pos})$ (30)
$=\frac{\partial}{\partial\widetilde{\bm{\mu}}}\log q(\widetilde{\bm{\beta}}\mid{\widetilde{\bm{\mu}}},\widetilde{\Sigma})-\frac{\partial}{\partial\widetilde{\bm{\mu}}}\log\phi(\widetilde{\bm{\beta}}\mid\bm{0},\widehat{\Sigma}_{1:k})-\sum_{i=1}^{n}\frac{\partial}{\partial\widetilde{\bm{\mu}}}\log L(y_{i};\bm{x}_{i}^{\top}\widetilde{\bm{\beta}})$ (31)
$=\bm{0}-\frac{\partial}{\partial\widetilde{\bm{\mu}}}\frac{-(\widetilde{\bm{\mu}}+\widetilde{\Sigma}_{1/2}\widetilde{\bm{\epsilon}})^{\top}\widehat{\Sigma}_{1:k}^{\dagger}(\widetilde{\bm{\mu}}+\widetilde{\Sigma}_{1/2}\widetilde{\bm{\epsilon}})}{2}-\sum_{i=1}^{n}\frac{\partial\bm{x}_{i}^{\top}\widetilde{\bm{\beta}}}{\partial\widetilde{\bm{\mu}}}\frac{\partial}{\partial\bm{x}_{i}^{\top}\widetilde{\bm{\beta}}}\log L(y_{i};\bm{x}_{i}^{\top}\widetilde{\bm{\beta}})$ (32)
$=\widehat{\Sigma}_{1:k}^{\dagger}(\widetilde{\bm{\mu}}+\widetilde{\Sigma}_{1/2}\widetilde{\bm{\epsilon}})-\sum_{i=1}^{n}(y_{i}-g(\bm{x}_{i}^{\top}\widetilde{\bm{\beta}}))\bm{x}_{i}$ (33)
$=\widehat{\Sigma}_{1:k}^{\dagger}\widetilde{\bm{\beta}}-\sum_{i=1}^{n}(y_{i}-g(\bm{x}_{i}^{\top}\widetilde{\bm{\beta}}))\bm{x}_{i},$ (34)
where $g(\cdot)$ is the logistic function. Similarly, the gradient with respect to $\widetilde{\Sigma}_{1/2}$ is given by
$\frac{\partial}{\partial\widetilde{\Sigma}_{1/2}}\mathrm{KL}(q(\widetilde{\bm{\beta}}\mid{\widetilde{\bm{\mu}}},\widetilde{\Sigma})\|p_{pos})$ (35)
$=\frac{\partial}{\partial\widetilde{\Sigma}_{1/2}}\log q(\widetilde{\bm{\beta}}\mid{\widetilde{\bm{\mu}}},\widetilde{\Sigma})-\frac{\partial}{\partial\widetilde{\Sigma}_{1/2}}\log\phi(\widetilde{\bm{\beta}}\mid\bm{0},\widehat{\Sigma}_{1:k})-\sum_{i=1}^{n}\frac{\partial}{\partial\widetilde{\Sigma}_{1/2}}\log L(y_{i};\bm{x}_{i}^{\top}\widetilde{\bm{\beta}})$ (36)
$=\frac{\partial}{\partial\widetilde{\Sigma}_{1/2}}\frac{-\log(|\widetilde{\Sigma}_{1/2}\widetilde{\Sigma}_{1/2}^{\top}|)-(\widetilde{\Sigma}_{1/2}\widetilde{\bm{\varepsilon}})^{\top}\widetilde{\Sigma}^{\dagger}(\widetilde{\Sigma}_{1/2}\widetilde{\bm{\varepsilon}})}{2}-\frac{\partial}{\partial\widetilde{\Sigma}_{1/2}}\frac{-\widetilde{\bm{\beta}}^{\top}\widehat{\Sigma}_{1:k}^{\dagger}\widetilde{\bm{\beta}}}{2}$ (37)
$\quad-\sum_{i=1}^{n}\frac{\partial\bm{x}_{i}^{\top}\widetilde{\bm{\beta}}}{\partial\widetilde{\Sigma}_{1/2}}\frac{\partial}{\partial\bm{x}_{i}^{\top}\widetilde{\bm{\beta}}}\log L(y_{i};\bm{x}_{i}^{\top}\widetilde{\bm{\beta}})$ (38)
$=-\widetilde{\Sigma}_{1/2}^{\dagger\top}+\frac{\partial}{\partial\widetilde{\Sigma}_{1/2}}\frac{(\widetilde{\bm{\mu}}+\widetilde{\Sigma}_{1/2}\widetilde{\bm{\epsilon}})^{\top}\widehat{\Sigma}_{1:k}^{\dagger}(\widetilde{\bm{\mu}}+\widetilde{\Sigma}_{1/2}\widetilde{\bm{\epsilon}})}{2}-\sum_{i=1}^{n}\frac{\partial\bm{x}_{i}^{\top}\widetilde{\bm{\beta}}}{\partial\widetilde{\Sigma}_{1/2}}\frac{\partial}{\partial\bm{x}_{i}^{\top}\widetilde{\bm{\beta}}}\log L(y_{i};\bm{x}_{i}^{\top}\widetilde{\bm{\beta}})$ (39)
$=-\widetilde{\Sigma}_{1/2}^{\dagger\top}+\frac{\partial}{\partial\widetilde{\Sigma}_{1/2}}\frac{(\widetilde{\Sigma}_{1/2}\widetilde{\bm{\epsilon}})^{\top}\widehat{\Sigma}_{1:k}^{\dagger}\widetilde{\Sigma}_{1/2}\widetilde{\bm{\epsilon}}}{2}+\frac{\partial}{\partial\widetilde{\Sigma}_{1/2}}\frac{2\widetilde{\bm{\mu}}^{\top}\widehat{\Sigma}_{1:k}^{\dagger}\widetilde{\Sigma}_{1/2}\widetilde{\bm{\epsilon}}}{2}$ (40)
$\quad-\sum_{i=1}^{n}\frac{\partial\bm{x}_{i}^{\top}\widetilde{\bm{\beta}}}{\partial\widetilde{\Sigma}_{1/2}}\frac{\partial}{\partial\bm{x}_{i}^{\top}\widetilde{\bm{\beta}}}\log L(y_{i};\bm{x}_{i}^{\top}\widetilde{\bm{\beta}})$ (41)
$=-\widetilde{\Sigma}_{1/2}^{\dagger\top}+\widehat{\Sigma}_{1:k}^{\dagger}\widetilde{\Sigma}_{1/2}\widetilde{\bm{\epsilon}}\widetilde{\bm{\epsilon}}^{\top}+\widehat{\Sigma}_{1:k}^{\dagger}\widetilde{\bm{\mu}}\widetilde{\bm{\epsilon}}^{\top}-\sum_{i=1}^{n}\bm{x}_{i}\widetilde{\bm{\epsilon}}^{\top}(y_{i}-g(\bm{x}_{i}^{\top}\widetilde{\bm{\beta}}))$ (42)
$=-\widetilde{\Sigma}_{1/2}^{\dagger\top}+\widehat{\Sigma}_{1:k}^{\dagger}\widetilde{\bm{\beta}}\widetilde{\bm{\epsilon}}^{\top}-\sum_{i=1}^{n}\bm{x}_{i}\widetilde{\bm{\epsilon}}^{\top}(y_{i}-g(\bm{x}_{i}^{\top}\widetilde{\bm{\beta}})),$ (43)
where $\widetilde{\Sigma}_{1/2}^{\dagger}$ denotes the pseudoinverse of $\widetilde{\Sigma}_{1/2}$.
The derived gradients can be used in conjunction with the Adam optimizer to update the variational parameters $\widetilde{\bm{\mu}}$ and $\widetilde{\Sigma}_{1/2}$. The optimization procedure iteratively updates the variational parameters until convergence or a maximum number of iterations is reached.

### v2

@@
 Abstract.
-The remarkable generalization performance of overparameterized models has challenged the conventional wisdom of statistical learning theory.
-While recent theoretical studies have shed light on this behavior in linear models or nonlinear classifiers, a comprehensive understanding of overparameterization in nonlinear regression models remains lacking.
-This paper explores the predictive properties of overparameterized nonlinear regression within the Bayesian framework, extending the methodology of adaptive prior based on the intrinsic spectral structure of the data.
-We establish posterior contraction for single-neuron models with Lipschitz continuous activation functions and for generalized linear models, demonstrating that our approach achieves consistent predictions in the overparameterized regime.
-Moreover, our Bayesian framework allows for uncertainty estimation of the predictions.
-The proposed method is validated through numerical simulations and a real data application, showcasing its ability to achieve accurate predictions and reliable uncertainty estimates.
-Our work advances the theoretical understanding of the blessing of overparameterization and offers a principled Bayesian approach for prediction in large nonlinear models.
+The remarkable generalization performance of large-scale models has been challenging the conventional wisdom of the statistical learning theory.
+Although recent theoretical studies have shed light on this behavior in linear models and nonlinear classifiers, a comprehensive understanding of overparameterization in nonlinear regression models is still lacking.
+This study explores the predictive properties of overparameterized nonlinear regression within the Bayesian framework, extending the methodology of the adaptive prior considering the intrinsic spectral structure of the data.
+Posterior contraction is established for generalized linear and single-neuron models with Lipschitz continuous activation functions, demonstrating the consistency in the predictions of the proposed approach.
+Moreover, the Bayesian framework enables uncertainty estimation of the predictions.
+The proposed method was validated via numerical simulations and a real data application, showing its ability to achieve accurate predictions and reliable uncertainty estimates.
+This work provides a theoretical understanding of the advantages of overparameterization and a principled Bayesian approach to large nonlinear models.
 1. Introduction
-In recent years, the field of machine learning has witnessed a paradigm shift in understanding the generalization performance.
-Traditional statistical theory emphasized the importance of controlling model complexity and avoiding overfitting by balancing the bias-variance tradeoff [55].
-However, the emergence of large-scale deep neural networks, which often have far more parameters than the number of training data, has challenged this conventional wisdom.
-These models have demonstrated the ability to achieve high accuracy on test data while completely interpolating the training data.
-This discrepancy between these empirical observations and traditional learning theory has necessitated a new explanation [61].
-A groundbreaking study by [5] provided a theoretical foundation in the overparameterized linear regression models, showing the non-asymptotic analysis of the generalization error of a minimum norm interpolator.
+In recent years, the field of machine learning has witnessed a paradigm shift in understanding generalization performance.
+Traditional statistical theory emphasizes the importance of controlling model complexity and avoids overfitting by balancing the bias-variance trade-off [56].
+However, with the emergence of large-scale deep neural networks, which often have considerably more parameters than the number of training data samples, conventional wisdom has been challenged.
+These models have demonstrated the ability to achieve high accuracy on test data by completely interpolating the training data.
+This discrepancy between empirical observations and traditional learning theory has necessitated a new explanation [62].
+A groundbreaking study by [6] provides a theoretical foundation for overparameterized linear regression models through a non-asymptotic analysis of the generalization error of a minimum norm interpolator.
 This discovery has reshaped our understanding of generalization performance in large statistical models, suggesting that overparameterization does not necessarily lead to poor prediction, even in simple models.
 Instead, the key to good performance lies in the model’s expressiveness and its ability to capture the essential structure of the data.
-Indeed, subsequent theoretical studies have demonstrated that overparameterization leads to accurate predictions in various statistical models, shedding light on the behavior of modern machine learning models.
-While the theoretical understanding of the goodness of overparameterization has advanced significantly in linear estimators, such as linear regression [5, 52] and kernel methods [31, 33], the analysis of nonlinear estimators remains limited.
-Although some studies have investigated this issue in nonlinear neural networks for classification problems [13, 26], a comprehensive understanding of why and under what conditions overparameterized nonlinear regression can achieve good generalization performance is still lacking.
+Recent theoretical studies have demonstrated that overparameterization leads to accurate predictions in various statistical models, shedding light on the behavior of modern machine learning models.
+While the theoretical understanding of the goodness of overparameterization has advanced significantly in linear estimation, such as in linear regression [6, 53] and kernel methods [32, 34], the analysis of nonlinear estimators remains limited.
+Although some studies have investigated this issue in nonlinear neural networks for classification problems [14, 27], a comprehensive understanding is lacking as to why and under what conditions overparameterized nonlinear regression can achieve good generalization performance.
 Hence, addressing this question is crucial for reinforcing the theoretical foundations of modern machine learning.
-In this paper, we explore the predictive properties of overparameterized nonlinear regression models within the Bayesian framework.
+In this paper, the predictive properties of overparameterized nonlinear regression models within the Bayesian framework are explored.
 Bayesian inference combines the prior distribution of parameters with a likelihood function to obtain the posterior distribution.
-[59] proposed a methodology that incorporates the structure of the data into the prior distribution while capturing the linear model structure as the likelihood, resulting in accurate predictions even in non-sparse high-dimensional models.
+[60] proposed a methodology that incorporates the structure of the data into the prior distribution, capturing the linear model structure as the likelihood, which results in accurate predictions even in non-sparse high-dimensional models.
 The key idea is to adaptively adjust the prior variance based on the intrinsic structure (spectra) of the data, which is found to be critical in prediction error convergence in the sense of posterior contraction.
-We extend this to nonlinear regression models, specifically single neuron models [60, 47] and generalized linear models (GLMs) [37], and theoretically analyze the prediction error of the resulting posterior distribution.
+We extend this to nonlinear regression models, specifically single-neuron models [61, 48] and generalized linear models (GLMs) [38] and theoretically analyze the prediction error of the resulting posterior distribution.
 The main contributions of this study are summarized as follows:
-  • For generalized linear models based on one-parameter exponential families, including logistic and Poisson regressions, we establish conditions under which the posterior distribution converges to the true parameter in terms of the predictive risk and provide an upper bound on the convergence rate (Theorem 1).
-This finding highlights the applicability of our approach to the regression of discrete response variables, such as those encountered in classification problems.
-  • We show that the proposed method generalizes well in single-neuron models with a Lipschitz continuous nonlinear activation function and Gaussian noise (Theorem 2).
-This result suggests that, even in more general models involving nonlinearity, overparameterization allows us to predict benignly under appropriate prior distributions.
-  • Our Bayesian approach not only provides a theoretical guarantee of the generalization performance of the overparametrized models but also estimates the uncertainty in the predictions.
-By sampling from the posterior distribution, we can construct a confidence level for each prediction, which is essential in decision-making processes.
+  • For generalized linear models based on one-parameter exponential families, including logistic and Poisson regressions, the conditions under which the posterior distribution converges to the true parameter are established in terms of predictive risk, providing an upper bound on the convergence rate (Theorem 1).
+This finding highlights the applicability of the proposed approach to discrete response-variable regression, such as those encountered in classification problems.
+  • The proposed method is demonstrated to generalize well to single-neuron models with a Lipschitz continuous nonlinear activation function and Gaussian noise (Theorem 2).
+This result suggests that even in more general models involving nonlinearity, overparameterization allows benign prediction under appropriate prior distributions.
+  • The proposed Bayesian approach not only provides a theoretical guarantee of the generalization performance of the overparametrized models but also estimates the uncertainty in the predictions.
+By sampling from the posterior distribution, a confidence level can be constructed for each prediction, which is essential in decision-making processes.
 1.1. Related work
-Bayesian GLM in high-dimension has been studied intensively.
-The first paper to study high-dimensional Bayesian GLM [18] achieved posterior concentration for the parameter without assuming any sparsity structure.
-However, it imposed the restricted growth condition on the parameter dimension, and this underparameterized setting differs from the present work.
-Subsequently, several studies have explored the GLM setting where the number of parameters may exceed the sample size [24, 34, 23].
-These works have demonstrated the asymptotic properties of the posterior distribution, but they assume sparsity, implying that the parameters are inherently low-dimensional.
+Bayesian GLM has been studied intensively for high dimensions.
+The first paper to study high-dimensional Bayesian GLM [19] realized posterior concentration for the parameter without assuming any sparsity structure by imposing a restricted growth condition on the parameter dimension.
+This underparameterized setting differs from the present work.
+Subsequently, several studies explored the GLM setting where the number of parameters may exceed the sample size [25, 35, 24].
+These works have demonstrated the asymptotic properties of the posterior distribution; however, they assume sparsity, implying that the parameters are inherently low-dimensional.
 The properties of Bayesian variable selection in GLMs have also been investigated in similar domains [9, 10, 43, 31, 23].
@@
 [6] demonstrated that the excess risk converges to zero in overparameterized linear models.
-This theory was widely extended in other linear estimators, such as regularized regression [52, 29, 58, 36, 11], kernel regression [31, 32], linear neural networks [10, 50], and regression for dependent data [40, 53].
-[13, 6, 26] have ventured beyond these linear frameworks to investigate the benign nature of overparameterization, but their focus is limited to classification problems.
-[47] tackled nonlinear regression and demonstrated the absence of benign overfitting, but their assumptions differ from the present work mainly in terms of distributional assumptions on the covariate.
+This theory has been widely extended to other linear estimators, such as regularized regression [53, 30, 59, 37, 12], kernel regression [32, 33], linear neural networks [11, 51], and regression for dependent data [41, 54].
+[14, 7, 27] ventured beyond these linear frameworks to investigate the benign nature of overparameterization, but their focus was limited to classification problems.
+[48] tackled nonlinear regression and demonstrated the absence of benign overfitting, but their assumptions differ from the present work mainly in terms of distribution of the covariate.
 1.2. Organization
 The remainder of this paper is structured as follows.
-In Section 2, we introduce the problem setting of one-parameter generalized linear models under the overparameterization assumption and provide a detailed description of our proposed methodology.
+In Section 2, the problem setting for one-parameter generalized linear models under the overparameterization assumption is introduced, providing a detailed description of the proposed methodology.
 Section 3 presents the theoretical results on posterior contraction for GLMs.
-Section 4 extends the results to a two-parameter case.
-We then conduct simulation studies in Section 5 to validate our theoretical findings and assess the uncertainty estimation.
-Section 6 showcases a real data application, comparing our approach with baseline methods in terms of classification performance.
-In Section 7, we conclude this paper by discussing the contribution and limitations of this work.
-The appendix provides the proof and discussion of the main theorems and the details of the computational algorithms used in our analysis.
+Section 4 extends the results to the two-parameter case.
+Section 5 reports on the simulation studies conducted to validate the theoretical findings and assess the uncertainty estimation.
+Section 6 showcases a real data application, comparing the proposed approach with baseline methods in terms of classification performance.
+Section 7 concludes the paper, discussing the contribution and limitations of this work.
+The proof and discussion of the main theorems and details of the computational algorithms used in the analysis are provided in the Appendix.
 1.3. Notation
@@
 For any vector $\bm{v}$ and symmetric matrix of the same dimension $A$, we define $\|\bm{v}\|_{A}=\sqrt{\bm{v}^{\top}A\bm{v}}$.
-For sequences $\{a_{n}\}_{n\in\mathbb{N}}$ and $\{b_{n}\}_{n\in\mathbb{N}}$, $a_{n}\gtrsim b_{n}$ denotes the existence of a constant $c>0$ such that $a_{n}\geq cb_{n}$ for all $n\geq\overline{n}$ with some finite $\overline{n}\in\mathbb{N}$, and $a_{n}\lesssim b_{n}$ denotes the opposite.
-$a_{n}=o(b_{n})$ indicates that $a_{n}/b_{n}\to 0$ as $n\to\infty$ whereas $a_{n}=\omega(b_{n})$ does $a_{n}/b_{n}\to\infty$ as $n\to\infty$.
-$a_{n}\gtrapprox b_{n}$ denotes that $a_{n}\gtrsim b_{n}(\log n)^{c}$ holds for every $c>0$, and $a_{n}\lessapprox b_{n}$ denotes the opposite.
+For sequences $\{a_{n}\}_{n\in\mathbb{N}}$ and $\{b_{n}\}_{n\in\mathbb{N}}$, $a_{n}\gtrsim b_{n}$ denotes the existence of a constant $c>0$ such that $a_{n}\geq cb_{n}$ for all $n\geq\overline{n}$ with some finite $\overline{n}\in\mathbb{N}$, and $a_{n}\lesssim b_{n}$ denotes the opposite; $a_{n}=o(b_{n})$ indicates that $a_{n}/b_{n}\to 0$ as $n\to\infty$, whereas $a_{n}=\omega(b_{n})$ indicates that $a_{n}/b_{n}\to\infty$ as $n\to\infty$; $a_{n}\gtrapprox b_{n}$ indicates that $a_{n}\gtrsim b_{n}(\log n)^{c}$ holds for every $c>0$, and $a_{n}\lessapprox b_{n}$ denotes the opposite.
 For natural numbers $a,b\in\mathbb{N}$ with $a<b$, we define $[a:b]:=\{a,a+1,a+2,...,b\}$.
@@
 2.1. Setting
-Consider a pair $(\bm{X},Y)$, where $\bm{X}$ is an $\mathbb{R}^{p}$-valued centered random covariate and $Y$ is an $\mathbb{R}$-valued univariate response. The GLM is defined as follows:
+Consider the pair $(\bm{X},Y)$, where $\bm{X}$ is an $\mathbb{R}^{p}$-valued centered random covariate, and $Y$ is an $\mathbb{R}$-valued univariate response. The GLM is defined as follows:
 $\mathbb{E}[Y\mid\bm{X}=\bm{x}]=g(\bm{x}^{\top}\bm{\beta}^{*}),$ (1)
 where $\bm{\beta}^{*}\in\mathbb{R}^{p}$ is the true unknown $p$-dimensional regression coefficient, and $g:\mathbb{R}\to\mathbb{R}$ is a known Lipschitz continuous function.
-In this setting, the response variable $Y$ follows a general distribution belonging to the one-parameter exponential family, with a density function given by:
+In this setting, the response variable $Y$ follows a general distribution belonging to the one-parameter exponential family with a density function given by
 $L(y;\mu_{x,\bm{\beta}})=\exp\{a(\mu_{x,\bm{\beta}})y+b(\mu_{x,\bm{\beta}})+c(y)\},$ (2)
-where $\mu_{x,\bm{\beta}}\in\mathbb{R}$ is the parameter, $a:\mathbb{R}\to\mathbb{R}$ and $b:\mathbb{R}\to\mathbb{R}$ are known continuously differentiable functions, with $a$ having a nonzero derivative.
+where $\mu_{x,\bm{\beta}}\in\mathbb{R}$ is the parameter; $a:\mathbb{R}\to\mathbb{R}$ and $b:\mathbb{R}\to\mathbb{R}$ are known continuously differentiable functions, with $a$ having a nonzero derivative.
 The function $c:\mathbb{R}\to\mathbb{R}$ is chosen such that $L(y;\mu_{x,\bm{\beta}})$ is a proper probability density.
 This formulation encompasses several well-known models: Normal linear models: $Y\mid\bm{X}$ follows a $g(\bm{x}^{\top}\bm{\beta}^{*})$-mean Gaussian distribution with fixed variance, and $g$ is the identity function; Logistic regression model: $Y\mid\bm{X}$ follows a Bernoulli distribution with parameter $g(\bm{x}^{\top}\bm{\beta}^{*})$, and $g$ is the logistic function; Poisson regression model: $Y\mid\bm{X}$ follows a Poisson distribution with parameter $g(\bm{x}^{\top}\bm{\beta}^{*})$, and $g$ is the exponential function.
-Note that we do not restrict the link function $g$ to canonical, providing flexibility in modeling the relationship between the linear predictor and the response variable.
+Note that the link function $g$ is not restricted to being canonical, thereby providing flexibility in modeling the relationship between the linear predictor and response variable.
 This study focuses on a non-sparse overparameterization setting, where $n$ independent and identically distributed copies of $(\bm{X},Y)$ are generated, and their realizations $\mathcal{D}:=\{(\bm{x}_{1},y_{1}),\ldots,(\bm{x}_{n},y_{n})\}$ are observed.
@@
 In the asymptotic setting, where $n$ diverges, $p$ possesses a larger order of divergence, i.e., $p=\omega(n)$.
-The primary goal of this study is to examine the generalization performance of a predictor trained on the observations $\mathcal{D}$ in the overparameterized domain.
+The primary objective of this study is to examine the generalization performance of a predictor trained on the observations $\mathcal{D}$ in the overparameterized domain.
 We adopt as the predictor the mean $g(\bm{x}^{\top}\bm{\beta})$, with the estimated parameter $\widehat{\bm{\beta}}$ plugged in.
@@
 2.2. Method
-The predictor construction follows a systematic approach: first, the data is divided into two sets; then, a prior distribution of $\bm{\beta}$ is developed based on the empirical spectral information of $\Sigma$ from one set; next, the likelihood is computed in the other set; finally, the posterior distribution of $\bm{\beta}$ is computed based on the prior and likelihood.
-To begin, we divide the dataset $\mathcal{D}$ into two equal parts, $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$, each containing $n/2$ observations.
-While the split ratio is arbitrary, it is crucial that it does not depend on the sample size $n$.
-This data splitting allows us to use $\mathcal{D}_{1}$ for constructing the prior distribution and $\mathcal{D}_{2}$ for evaluating the likelihood function, ensuring that the data used in these two steps are independent.
+Predictor construction follows a systematic approach: first, the data are divided into two sets; then, a prior distribution of $\bm{\beta}$ is developed based on the empirical spectral information of $\Sigma$ from one set; next, the likelihood is computed in the other set; finally, the posterior distribution of $\bm{\beta}$ is computed based on the prior and likelihood.
+Initially, we divide the dataset $\mathcal{D}$ into two equal parts, $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$, each containing $n/2$ observations.
+Although the split ratio is arbitrary, it should not depend on the sample size $n$.
+The data splitting allows the use of $\mathcal{D}_{1}$ for constructing the prior distribution and $\mathcal{D}_{2}$ for evaluating the likelihood function, ensuring that the data used in these two steps are independent.
 Next, we consider the empirical covariance matrix of the covariates $\bm{X}$, defined as:
@@
 where $\bar{\bm{x}}\in\mathbb{R}^{p}$ is the sample mean of the covariates from $\mathcal{D}_{1}$.
-The factor $2/n$ ensures that $\widehat{\Sigma}$ is an unbiased estimator of the population covariance matrix.
-We then perform a spectral decomposition of $\widehat{\Sigma}$:
+The factor $2/n$ ensures that $\widehat{\Sigma}$ is an unbiased estimator of the population covariance matrix, whereby a spectral decomposition of $\widehat{\Sigma}$ is performed:
 $\widehat{\Sigma}=\sum_{j=1}^{p}\widehat{\lambda}_{j}\widehat{\bm{v}}_{j}\widehat{\bm{v}}_{j}^{\top},$ (4)
@@
 This approximation captures the most significant $k$ eigenvalues and eigenvectors of $\widehat{\Sigma}$, effectively reducing the dimensionality of the covariate space while retaining the most informative directions of variation.
-The choice of $k$ is a tuning parameter; however, to manage computational cost, we fix it in an interval $[L_{\kappa}:U_{\kappa}]$ such that it satisfies Assumption 2.
+In choosing $k$ as a tuning parameter to manage the computational cost, it has to be fixed to an interval of $[L_{\kappa}:U_{\kappa}]$ so that it satisfies Assumption 2.
 We then employ the following prior distribution for $\bm{\beta}$.
@@
 Second, unlike [43, 24, 55], we impose a non-sparse prior distribution, even though some parameters may be redundant.
-This choice is motivated by the desire to retain all potentially relevant covariates in the data and to avoid the challenges associated with sparse priors, such as the need to specify sparsity-inducing hyperparameters.
+This choice is motivated by the desire to retain all potentially relevant covariates in the data and avoid the challenges associated with sparse priors, such as the need to specify sparsity-inducing hyperparameters.
 Next, we consider the posterior distribution, which combines the prior distribution and the likelihood function using Bayes’ Theorem.
@@
 $\Pi\left(\{\bm{\beta}:\bm{\beta}\in B\}\mid\mathcal{D},k\right)=\int_{B}\pi(\bm{\beta}\mid\mathcal{D},k)\mathrm{d}\bm{\beta}.$ (8)
-The Bayesian framework allows for the quantification of events of interest, such as the probability of predictions being close to the true values, through the probability measures defined in this manner.
-Moreover, as will be demonstrated later, the distributional information provided by the posterior is valuable in practice.
-Note that we discard the information about the response variable $y_{i}$ in $\mathcal{D}_{1}$ since this subset was only used to compute the sample covariance matrix of the covariates.
-If dropping this information seems inefficient, one can consider switching the roles of $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ to construct a new estimator (posterior distribution), and averaging the two estimators in a manner similar to the sample splitting scheme [7].
-The theoretical results presented later are preserved for this averaging.
+The Bayesian framework allows the quantification of events of interest, such as the probability of predictions being close to the true values through probability measures defined in this manner.
+Moreover, as demonstrated later, the distributional information provided by the posterior is valuable in practice.
+Note that the information about the response variable $y_{i}$ in $\mathcal{D}_{1}$ is discarded since this subset is only used to compute the sample covariance matrix of the covariates.
+If excluding this information seems inefficient, the roles of $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ can be switched to construct a new estimator (posterior distribution), and the two estimators can be averaged in a manner similar to the sample splitting scheme [8].
+The theoretical results presented next preserve this averaging.
 3. Main Results
-In this section, we present the main theoretical results.
-We begin by introducing the key assumptions that underlie our analysis (§3.1).
+In this section, the main theoretical results are presented.
+Initially, the key assumptions that underlie our analysis (§3.1) are introduced.
 These assumptions concern the spectral properties of the covariance matrix and the concentration of the empirical covariance matrix around its population counterpart.
-We then proceed to establish the posterior contraction for the GLM (§3.2).
-Our theoretical results offer characterizations of the posterior contraction, elucidating the intricate interplay between the dimensionality, sample size, and the eigenvalue decay of the covariance matrix.
+Subsequently, the posterior contraction for the GLM is established (§3.2).
+The theoretical results characterize the posterior contraction, elucidating the intricate interplay between dimensionality, sample size, and eigenvalue decay of the covariance matrix.
 3.1. Assumptions
-We begin by focusing on the boundedness of the operator norm of $\Sigma$ and the concentration of the empirical covariance matrix $\widehat{\Sigma}$ around $\Sigma$.
+The initial focus is on the boundedness of the operator norm of $\Sigma$ and concentration of the empirical covariance matrix $\widehat{\Sigma}$ around $\Sigma$.
 Assumption 1.
@@
 $\mathbb{P}\left(\frac{\|\widehat{\Sigma}-\Sigma\|_{\mathrm{op}}}{\|\Sigma\|_{\mathrm{op}}}\leq\rho_{n}\right)\geq 1-\ell_{n}.$ (9)
-This assumption ensures that the empirical covariance matrix concentrates well around the population covariance matrix.
-The sequence $\rho_{n}$ quantifies the rate of this concentration, while $\ell_{n}$ controls the probability of the concentration event.
-It is worth noting that this assumption is mild and includes not only the sub-Gaussian distributions commonly imposed in high-dimensional statistics but also log-concave distributions, which encompass most unimodal parametric distributions [56, 57] and allow for heavy-tailed distributions such as the Laplace distribution.
-In contrast, other related literature has imposed more restrictive conditions such as sub-Gaussianity [5] or strong log-concavity [13].
+This assumption ensures that the empirical covariance matrix concentrates around the population covariance matrix.
+The sequence $\rho_{n}$ quantifies the rate of this concentration, whereas $\ell_{n}$ controls the probability of the concentration event.
+Notably, this assumption is mild and includes not only the sub-Gaussian distributions commonly imposed in high-dimensional statistics but also log-concave distributions, which encompass most unimodal parametric distributions [57, 58] that allow for heavy-tailed distributions such as the Laplace distribution.
+In contrast, other related literature impose more restrictive conditions such as sub-Gaussianity [6] or strong log-concavity [14].
 In a high-dimensional setting, where many covariates are incorporated, the mildness of the distributional assumption is particularly important, as it allows for a wider range of applicable scenarios.
-Our next assumption concerns the behavior of the eigenvalues of $\Sigma$.
+The next assumption concerns the behavior of the eigenvalues of $\Sigma$.
 Assumption 2.
-Given a covariance matrix $\Sigma$ satisfying (9), with eigenvalues $\{\lambda_{k}\}_{k=1}^{p}$, the sequence $\{\rho_{n}\}_{n\in\mathbb{N}}$ from Assumption 1, and a sequence of bounds $(L_{\kappa},U_{\kappa})$, there exists a sequence $\{\varepsilon_{n}\}_{n\in\mathbb{N}}$ of positive reals such that the following hold as $n\to\infty$:
+Given a covariance matrix $\Sigma$ satisfying (9), with eigenvalues $\{\lambda_{k}\}_{k=1}^{p}$, sequence $\{\rho_{n}\}_{n\in\mathbb{N}}$ from Assumption 1, and a sequence of bounds $(L_{\kappa},U_{\kappa})$, there exists a sequence $\{\varepsilon_{n}\}_{n\in\mathbb{N}}$ of positive reals such that the following hold as $n\to\infty$:
 $\mathrm{(i)}$ $\quad\varepsilon_{n}\to 0,\quad n^{-1/2}\lesssim\rho_{n}\lesssim\varepsilon_{n}^{2},\qquad\mathrm{(ii)}\quad(\rho_{n}+\lambda_{k})\lambda_{k}\lessapprox\frac{\varepsilon_{n}^{2}}{k},\quad\forall k\in[L{\kappa}:U_{\kappa}],$ (10)
@@
 Intuitively, conditions (i) and (ii) control the rate at which the bulk of the eigenvalues (those indexed from $L_{\kappa}$ to $U_{\kappa}$) are allowed to decay moderately.
-Condition (iii) requires that the eigenvalues beyond the bulk (those indexed larger than $U_{\kappa}$) are sufficiently small, while condition (iv) constrains the size of $U_{\kappa}$ relative to the sample size $n$ and the rate $\varepsilon_{n}$.
+Condition (iii) requires that the eigenvalues beyond the bulk (those indexed larger than $U_{\kappa}$) are sufficiently small, whereas condition (iv) constrains the size of $U_{\kappa}$ relative to sample size $n$ and rate $\varepsilon_{n}$.
 These conditions are crucial for deriving the posterior contraction rates in Theorem 1.
-Finally, we assume a certain local Lipschitz condition of the link function $g$ in terms of the Kullback-Leibler (KL) divergence and the variational form between the corresponding probability distributions.
+Finally, we assume a certain local Lipschitz condition of the link function $g$ in terms of the Kullback-Leibler (KL) divergence and variational form between the corresponding probability distributions.
 Assumption 3.
 For any pair $\bm{\beta},\bm{\beta}^{\prime}\in\mathbb{R}^{p}$, the GLM with link function $g$ satisfies
-$K(P,P^{\prime}),~{}V(P,P^{\prime})\lesssim\|\bm{\beta}-\bm{\beta}^{\prime}\|_{\Sigma}^{2},$
+$K(P,P^{\prime}),V(P,P^{\prime})\lesssim\|\bm{\beta}-\bm{\beta}^{\prime}\|_{\Sigma}^{2},$
 where $P$ and $P^{\prime}$ are the probability distributions corresponding to the densities $L(y;g(\bm{x}^{\top}\bm{\beta}))$ and $L(y;g(\bm{x}^{\top}\bm{\beta}^{\prime}))$, respectively.
-This assumption limits the amount of variation in a distribution when different parameters are used, meaning that it controls the complexity of the model.
-This assumption is satisfied by several common GLMs, as illustrated in the following examples.
+This assumption limits the amount of variation in a distribution when different parameters are used, to control the complexity of the model and is satisfied by several common GLMs, as illustrated in the following examples.
 Example 1.
@@
 Theorem 1.
-Consider the GLM (2) with covariate distribution satisfying Assumptions 1-3, and a prior distribution on $\bm{\beta}$ is (6).
+Consider the GLM (2) with covariate distribution satisfying Assumptions 1-3, and the prior distribution on $\bm{\beta}$ is (6).
 Then, for a sufficiently large constant $M$, the posterior distribution contracts around the true parameter $\bm{\beta}^{*}$ as
-[…]
+$\Pi\Big{(}{\bm{\beta}:\|g(\bm{X}^{\top}\bm{\beta})-g(\bm{X}^{\top}\bm{\beta}^{*})\|_{P_{\bm{X}},2}>M\varepsilon_{n}}\mid,\mathcal{D}\Big{)}\overset{P^{*}}{\longrightarrow}0\quad\mbox{as}\quad n\to\infty.$
 Furthermore, if the eigenvalues of $\Sigma$ satisfy the additional condition $(nk)^{-1}\gtrsim(\rho_{n}+\lambda_{k})\lambda_{k}$ for all $k\in[L_{\kappa}:U_{\kappa}]$, then the contraction rate can be explicitly characterized as
 $\varepsilon_{n}=O\left(U_{\kappa}\lambda_{L_{\kappa}}^{1/2}(\lambda_{L_{\kappa}}+\rho_{n})^{1/2}\right).$ (12)
-This theorem shows that under suitable conditions on the covariance matrix $\Sigma$ and the prior distribution, the posterior distribution of the predictor concentrates around the oracle predictor at a rate $\varepsilon_{n}$ with respect to the $L_{2}(P_{\bm{X}})$ distance.
-The rate $\varepsilon_{n}$ depends on the interplay between the sample size $n$, the concentration rate $\rho_{n}$ of the empirical covariance matrix, and the eigenvalues of $\Sigma$ in the bulk region.
+This theorem shows that under conditions suitable for the covariance matrix $\Sigma$ and prior distribution, the posterior distribution of the predictor concentrates around the oracle predictor at a rate $\varepsilon_{n}$ with respect to the $L_{2}(P_{\bm{X}})$ distance.
+The rate $\varepsilon_{n}$ depends on the interplay between sample size $n$, concentration rate $\rho_{n}$ of the empirical covariance matrix, and eigenvalues of $\Sigma$ in the bulk region.
 3.2.1. Consistency as a distribution
-While Theorem 1 derives posterior consistency of the predictor with respect to the $L_{2}$ distance, it also holds for other commonly used distances between probability distributions, such as the total variation distance, the Hellinger distance, and the KL divergence in several cases, including a list of distributions in Example 1.
-The contraction in these divergences is particularly useful for assessing the quality of the posterior distribution in terms of its ability to approximate the true data-generating distribution.
-A detailed discussion of these results is provided in the appendix.
+While Theorem 1 derives the posterior consistency of the predictor with respect to the $L_{2}$ distance, it also holds for other commonly used distances between probability distributions, such as the total variation distance, the Hellinger distance, and the KL divergence in several cases, including a list of distributions in Example 1.
+The contraction in these divergences is particularly useful for assessing the quality of the posterior distribution in approximating the true data-generating distribution.
 3.2.2. Consistency of Point Estimators
@@
 $\widehat{\bm{\beta}}_{\text{MAP}}=\operatornamewithlimits{argmax}_{\bm{\beta}\in\mathbb{R}^{p}}\Pi(\bm{\beta}\mid\mathcal{D}).$
-Under the same assumptions as in Theorem 1, it can be shown that the MAP estimator is consistent with respect to the $L_{2}(P_{\bm{X}})$ norm:
+Under the same assumptions as in Theorem 1, the MAP estimator can be shown to be consistent with respect to the $L_{2}(P_{\bm{X}})$ norm:
 $\|g(\bm{X}^{\top}\widehat{\bm{\beta}}_{\text{MAP}})-g(\bm{X}^{\top}\bm{\beta}^{*})\|_{P_{\bm{X}},2}\overset{P^{*}}{\longrightarrow}0,\quad\mbox{as}~{}n\to\infty.$
@@
 […]
-Under the assumptions of Theorem 1 and the additional condition that posterior consistency with respect to the total variation distance or the Hellinger distance holds (as discussed in §3.2.1), the posterior mean is also consistent:
+Under the assumptions of Theorem 1 and the additional condition that posterior consistency with respect to the total variation distance or Hellinger distance holds (as discussed in §3.2.1), the posterior mean is also consistent:
 $\|g(\bm{X}^{\top}\widehat{\bm{\beta}}_{\text{PM}})-g(\bm{X}^{\top}\bm{\beta}^{*})\|_{P_{\bm{X}},2}\overset{P^{*}}{\longrightarrow}0,\quad\mbox{as}~{}n\to\infty.$ (14)
-As the posterior mean considers the weights of the entire distribution (as it is defined by integral), its consistency requires the additional condition that the posterior distribution is not ill-behaved.
-We defer the further discussion to the appendix.
+As the posterior mean considers the weights of the entire distribution (defined by integral), its consistency requires the additional condition that the posterior distribution is not ill-behaved.
+Further discussion is deferred to the Appendix.
 4. Extension to Two-Parameter Gaussian Case
-In this section, we extend our analysis to the setting where the GLM has a Gaussian distribution for the measurement error.
-Specifically, we consider the following model, referred to as the single neuron model [47, 60]:
+In this section, the analysis is extended to the setting of GLM with a Gaussian distribution for the measurement error. Specifically, we consider the single-neuron model [48, 61]:
 $Y=g(\bm{X}^{\top}\bm{\beta})+\varepsilon;\quad\varepsilon\sim N(0,\sigma^{2}),$ (15)
@@
 For instance, if $g$ is a censored link function such as the ReLU, tanh, or softplus, then (15) represents a censored regression model.
-Here, we specify the prior distributions for the parameters $\bm{\beta}$ and $\sigma^{2}$.
-For $\bm{\beta}$, we continue to use the prior distribution (6).
-Regarding the variance parameter, we employ as the prior distribution the inverse Gaussian distribution $\Pi_{\varsigma^{2}}(\cdot)$ with the density of the form:
+Here, after specifying the prior distributions for the parameters $\bm{\beta}$ and $\sigma^{2}$, the prior distribution (6) for $\bm{\beta}$ is used.
+Regarding the variance parameter, we employ as the prior distribution, the inverse Gaussian distribution $\Pi_{\varsigma^{2}}(\cdot)$ with density of the form:
 $\pi_{\varsigma^{2}}(\sigma^{2})=\sqrt{\frac{\xi}{2\pi\sigma^{6}}}\exp\left(\frac{-\xi(\sigma^{2}-\eta)}{2\eta^{2}\sigma^{2}}\right),$ (16)
-where $\eta>0$ and $\xi>0$ denote the mean and shape parameters, respectively.
-We allow for arbitrary assignment of hyperparameters.
+where $\eta>0$ and $\xi>0$ denote the mean and shape parameters, respectively, and the hyperparameters are arbitrarily assigned.
 An important property of the inverse Gaussian prior is that the tails on both sides decrease rapidly, which is critical for posterior contraction.
@@
 Theorem 2.
-Consider the regression model (16) with covariate distribution satisfying Assumptions 1-3, and a prior distribution on $(\bm{\beta},\sigma^{2})$ satisfying (6) and (17).
+Consider the regression model (15) with covariate distribution satisfying Assumptions 1-3 and a prior distribution on $(\bm{\beta},\sigma^{2})$ satisfying (6) and (16).
 Then, for a sufficiently large constant $M$, the following convergences hold as $n\to\infty$:
@@
 $\varepsilon_{n}=O\left(U_{\kappa}\lambda_{L_{\kappa}}^{1/2}(\lambda_{L_{\kappa}}+\rho_{n})^{1/2}\right).$ (19)
-This theorem provides the posterior contraction for both the predictor $g(\bm{X}^{\top}\bm{\beta})$ and the error variance $\sigma^{2}$ in the two-parameter Gaussian model.
-The contraction rate for the predictor is with respect to the $L_{2}(P_{\bm{X}})$ distance, while the rate for $\sigma^{2}$ is with respect to the absolute difference.
-The extension to the two-parameter Gaussian model is particularly relevant for applications where the error variance is unknown and needs to be estimated from the data.
+This theorem provides the posterior contraction for both predictor $g(\bm{X}^{\top}\bm{\beta})$ and error variance $\sigma^{2}$ in the two-parameter Gaussian model.
+The contraction of the predictor is with respect to the $L_{2}(P_{\bm{X}})$ distance, while that of $\sigma^{2}$ is with respect to the absolute difference.
+The extension of the two-parameter Gaussian model is particularly relevant to applications where the error variance is unknown and needs to be estimated from the data.
 By considering the joint posterior contraction rates for $(\bm{\beta},\sigma^{2})$, Theorem 2 presents a theoretical justification for Bayesian inference in such models and offers guidance on the choice of prior distributions that can adapt to the unknown error variance.
 5. Simulation
-This section confirms by simulation the statement of the theorem presented in §3 and §4.
+This section confirms by simulation the statement of the theorem presented in §3 and §4 through simulations.
+The posterior predictive distribution of the proposed method is computed using variational inference, and the detailed algorithm is deferred to the Appendix.
 5.1. Logistic Regression
-Here, we consider a binary classification scenario where the response variable $Y_{i}$ takes values in $\{0,1\}$. The response variables are generated according to the following model:
-$Y_{i}=\mathrm{Ber}(g_{ls}(\bm{X}_{i}^{\top}\bm{\beta}));\quad i=1,2,...,n,$
+Here, we consider a binary classification scenario where the response variable $Y_{i}$ takes values in $\{0,1\}$. The data are generated according to the following model:
+$Y_{i}=\mathrm{Ber}(g_{ls}(\bm{X}_{i}^{\top}\bm{\beta}));\quad i=1,2,...,n,$ (20)
 where $g_{ls}(r):=1/(1+e^{-r})$ is the logistic function.
-We set $\bm{\beta}\sim N(\bm{0},I_{p})$ and $p=n^{4/3}$, ensuring an overparameterized setting.
-Regarding the covariate, we consider two settings: (A) Gaussian covariate, that is, $\bm{X}_{i}\sim N(\bm{0},\Sigma)$; (B) Laplacian convariate, that is, $\bm{X}_{i}\sim L(\bm{0},\Sigma)$, where the eigenvalues of $\Sigma$ follow $\lambda_{j}(\Sigma)=10\exp(-j/8)+n\exp\{-\sqrt{n}\}/p$.
-We investigate the behavior of the proposed method as the sample size increases by varying the training data size $n$ from $50$ to $500$.
-For each sample size, we generate a test dataset of $1000$ instances to evaluate the performance of the trained classifier.
-To evaluate the performance of the Bayesian predictions, we employ four widely used metrics: 0-1 loss, which quantifies the misclassification rate; Area Under the ROC Curve (AUC), which assesses the discriminative ability of the classifier; Unconfident Misclassifications rate (UM), which measures the proportion of unconfident predictions among misclassified instances; and Confident Correct predictions rate (CC), which calculates the proportion of accurate predictions among confident ones.
-We refer to a confident prediction as one in which the $95\%$ interval does not cross the threshold $(0.5)$, while an unconfident one does.
+Specifically, for $n=50,100,\ldots,500$, to ensure an overparameterized setting, $\bm{\beta}\sim N(\bm{0},I_{p})$ are generated with $p=n^{4/3}$.
+Two settings are considered for the covariate: (A) Gaussian covariate, $\bm{X}_{i}\sim N(\bm{0},\Sigma)$; (B) Laplacian covariate, $\bm{X}_{i}\sim L(\bm{0},\Sigma)$, where the eigenvalues of $\Sigma$ follow $\lambda_{j}(\Sigma)=10\exp(-j/8)+n\exp\{-\sqrt{n}\}/p$.
+Then, each $Y_{i}$ corresponding to $X_{i}$ and $\beta$ is sampled from (20).
+For every $n$, $20$ different datasets were repeatedly generated through the above procedure and each was analyzed to determine the behavior of the proposed method in response to an increase in sample size.
+In each case, a test dataset of $1000$ instances was generated, to evaluate the performance of the trained classifier.
+The performance of the Bayesian predictions was evaluated, employing four widely used metrics: 0-1 loss, which quantifies the misclassification rate; area under the ROC curve (AUC), which assesses the discriminative ability of the classifier; rate of unconfident misclassifications (UM), which measures the proportion of unconfident predictions among misclassified instances; and confident correct (CC) predictions rate, which calculates the proportion of accurate predictions among confident ones.
+We refer to a confident prediction as one whose $95\%$ predictive interval does not cross the threshold $(0.5)$, whereas an unconfident one does.
 The latter two metrics are helpful when building systems that defer action when there is no confidence in the prediction.
+Additionally, principal component regression (PCR) [5] was performed as a baseline method, with the number of principal components selected by 5-fold cross-validation from $\{1,2,\ldots,30\}$.
 Figure 1.
-Change of 0-1 loss, AUC (area under the ROC curve), UM (proportion of unconfident predictions among misclassifications), and CC (proportion of accurate predictions among confident ones) as the sample size grows in the logistic regression with the Gaussian covariate case (left) and the Laplace covariate case (right).
-Figure 1 illustrates the results of the experiment.
-As the sample size increases, the 0-1 loss exhibits a consistent decrease in both settings, indicating an improvement in classification accuracy.
+Change of 0-1 loss, AUC (area under the ROC curve), UM (proportion of unconfident predictions among misclassifications), and CC (proportion of accurate predictions among confident ones) of the proposed method as sample size increases for the cases of logistic regression with Gaussian covariate (left) and Laplace covariate (right).
+The gray lines represent the results of the principal component regression.
+The shaded bands represent the 90% interval.
+Figure 1 illustrates the results, where the reported values are averaged over the $20$ datasets.
+As sample size increases, the 0-1 loss exhibits a consistent decrease in both settings, indicating an improvement in classification accuracy.
 This trend persists even in the overparameterized setting, aligning with the theoretical findings.
-The AUC metric also shows a steady increase with the sample size, suggesting an enhanced ability to discriminate between the two classes.
+The AUC metric also shows a steady increase with sample size, suggesting an enhanced ability to discriminate between the two classes.
+Our method outperformed PCR, and these results underscore the importance of overparameterization for prediction.
 The UM and CC metrics provide insights into the uncertainty estimation of the proposed method.
 UM consistently presents a high value.
-This indicates that the method is able to warn us of a lack of confidence when misclassifying.
-The CC also presents a similarly high value, which means that confident predictions are trustworthy.
-These observations suggest that the proposed method effectively captures the uncertainty associated with predictions and leverages it to make more informed decisions.
+This indicates that the method can warn us of a lack of confidence when misclassifying.
+The CC also presents a similarly high value, suggesting that confident predictions are trustworthy.
+These observations suggest that the proposed method effectively captures the uncertainty associated with predictions, thereby leveraging it to make more informed decisions.
 5.2. Nonlinear regression with Gaussian noise
-In this subsection, we consider a scenario where the true response takes values in the positive real line $\mathbb{R}_{+}$. Specifically, the response variables are generated as follows:
-$Y_{i}=N(g_{sp}(\bm{X}_{i}^{\top}\bm{\beta}),1);\quad i=1,2,...,n,$
-where $\bm{\beta}\sim N(\bm{0},I_{p})$, $p=n^{4/3}$, and $g_{sp}(r):=\log(1+e^{r})$ is the softplus function.
-The distributions of the covariate vectors are set to follow the same patterns as in the previous experiment.
-We examine how the performance of Bayesian predictions changes as the data size increases by varying the training data size $n$ from $50$ to $500$.
-For a newly generated test dataset $\{(\bm{x}_{i},g_{sp}(\bm{x}_{i}^{\top}\bm{\beta}))\}_{i=1}^{1000}$, we assess the accuracy of Bayesian predictions using three widely adopted indicators: Root Mean Squared Error (RMSE), measuring the average deviation between the predicted values and the true values; Coverage Probability (CP), evaluating the reliability of the $95\%$ prediction intervals by calculating the proportion of true values falling within the intervals; and Average Length (AL), quantifying the width of the $95\%$ prediction intervals, providing insight into the precision of the predictions.
+In this subsection, the scenario of true responses taking values along the positive real line $\mathbb{R}_{+}$ is considered. Specifically, the data are generated as follows:
+$Y_{i}=N(g_{sp}(\bm{X}_{i}^{\top}\bm{\beta}),1);\quad i=1,2,...,n,$ (21)
+where $g_{sp}(r):=\log(1+e^{r})$ is the softplus function.
+For $n=50,100,\ldots,500$, we set $p=n^{4/3}$ and sample the coefficient $\bm{\beta}\in\mathbb{R}^{p}$ and covariate vectors $X_{i}\in\mathbb{R}^{p}$ in the same manner as in the previous experiment.
+Then, the scalar response is generated by (21).
+For every $n$, $20$ distinct datasets were generated using the above procedure and, on each one, Bayesian prediction performance changes were examined with an increase in training data size.
+For a newly generated test dataset $\{(\bm{x}_{i},g_{sp}(\bm{x}_{i}^{\top}\bm{\beta}))\}_{i=1}^{1000}$, the accuracy of Bayesian predictions was assessed using three widely adopted indicators: root-mean-squared error (RMSE), measuring the average deviation between predicted and true values; coverage probability (CP), evaluating the reliability of the $95\%$ prediction intervals by calculating the proportion of true values falling within the confidence intervals; and average length (AL), quantifying the width of the $95\%$ prediction intervals, providing insight into the precision of the predictions.
+Moreover, PCR was implemented, selecting the number of principal components using 5-fold cross-validation from $\{1,2,\ldots,30\}$.
 Figure 2.
-Change of RMSE (root mean squared error), CP (coverage probability of $95$% prediction interval), and AL (average length of $95$% prediction interval) as the sample size grows in nonlinear Gaussian regression with the Gaussian covariate case (left) and the Laplace covariate case (right).
-The results are shown in Figure 2.
-First, the RMSE decreases monotonically as the sample size increases.
-Even when $p$ becomes large and the setting becomes clearly overparameterized, the fact that the accuracy continues to improve supports the theoretical results experimentally.
-Next, we focus on CP and AL.
-These are indicators related to prediction intervals, with CP representing the reliability of the prediction interval and AL representing its width.
-From the figure, CP is not good when $n=100$, but after that, it stabilizes at a high value.
-AL is also slightly increasing, but not too much, implying that the interval is not trivial.
-(Regardless of $n$, the mean of $\{g_{sp}(\bm{x}_{i}^{\top}\bm{\beta})\}_{i=1}^{1000}$ in the test data is around $3.5$ and the variance is around $5$.) These results suggest that the proposed method is able to properly quantify the uncertainty in constructing prediction intervals.
+Change of RMSE (root mean squared error), CP (coverage probability of $95$% prediction interval), and AL (average length of $95$% prediction interval) as sample size increases in nonlinear Gaussian regression with the Gaussian covariate case (left) and Laplace covariate case (right).
+The results, averaged over $20$ datasets, are shown in Figure 2.
+First, the RMSE decreases monotonically as sample size increases.
+Even when $p$ is large and the setting is clearly overparameterized, the fact that accuracy continues to improve supports the theoretical results experimentally.
+Also, our method excels PCR, providing advantages of our method for prediction problems.
+Next, we focus on CP and AL, which are indicators related to prediction intervals, with CP representing the reliability of the prediction interval and AL representing its width.
+In the figure, the coverage is not sufficient when $n=100$; however, after that, CP stabilizes at a high value.
+AL also slightly increases but not by much, implying that the interval is not trivial.
+(Regardless of $n$, the mean of $\{g_{sp}(\bm{x}_{i}^{\top}\bm{\beta})\}_{i=1}^{1000}$ in the test data is around $3.5$, and the variance is around $5$.) These results suggest that the proposed method properly quantifies the uncertainty in constructing prediction intervals.
 6. Data Analysis
@@
 The dataset consists of mass spectrometric measurements from cancer patients and healthy individuals, with the goal of distinguishing between the two groups based on these measurements.
-The ARCENE dataset, accessible through the UC Irvine Machine Learning Repository [16], was constructed by merging three separate datasets: National Cancer Institute (NCI) ovarian cancer data; NCI prostate cancer data; and Eastern Virginia Medical School (EVMS) prostate cancer data.
+The ARCENE dataset, accessible through the UC Irvine Machine Learning Repository [17], was constructed by merging three separate datasets: National Cancer Institute (NCI) ovarian cancer data, NCI prostate cancer data, and Eastern Virginia Medical School (EVMS) prostate cancer data.
 In total, the public dataset contains $200$ samples, comprising $88$ cancer samples and $112$ normal samples.
-These data are categorized as cancer samples and control samples, with $10000$ continuous covariates, with 7000 dimensions corresponding to real mass-spectrometry peaks and $3000$ dimensions representing random noise variables (probes).
+These data are categorized as cancer and control samples, with $10000$ continuous covariates, with 7000 dimensions corresponding to real mass-spectrometry peaks and $3000$ dimensions representing random noise variables (probes).
 The primary challenges posed by the dataset lie in its high dimensionality and the presence of overlapping classes.
-The goal here is to discriminate between cancer and normal samples.
+The objective here is to discriminate between cancerous and normal samples.
 6.2. Experimental Setting
-We compared the performance of five distinct methods.
-  • Logistic Regression with the effective spectral (ES) prior distribution: The proposed Bayesian methodology, as described in Section 2.2, was applied to the logistic regression model to make predictions on unseen data.
-  • Logistic Regression with the Horseshoe (HS) prior: To handle the high-dimensionality of covariates, we implemented logistic regression with the Horseshoe prior [12].
-The implementation was carried out using the ”horseshoenlm” package in R [38].
+The performance of the six distinct methods were compared.
+  • Logistic Regression with effective spectral (ES) prior distribution: The proposed Bayesian methodology, as described in Section 2.2, was applied to the logistic regression model to make predictions on unseen data.
+  • Logistic Regression with the horseshoe (HS) prior: To handle the high-dimensionality of covariates, logistic regression was implemented using the Horseshoe prior [13].
+The ”horseshoenlm” package in R [39] was used for the implementation.
   • Logistic Regression with the normal (NM) prior: We investigated the performance of logistic regression with a standard normal prior distribution, which is the most commonly used prior and serves as a ridge-like regularization [16].
-This prior allows for flexibility in modeling the coefficient distribution and can be easily implemented using the Pyro probabilistic programming language [3].
-  • Logistic Regression with the Laplace (LP) prior: We also explored the performance of logistic regression with a standard Laplace prior distribution, which promotes sparsity in the coefficient estimates as a Bayesian Lasso [43].
-This prior has been shown to be effective in high-dimensional settings and can be efficiently implemented using Pyro, leveraging its support for non-Gaussian prior distributions [3].
-  • LightGBM: As a baseline method for table data prediction tasks, we implemented LightGBM [27], an efficient and scalable gradient boosting framework.
-We adopted the ”lightgbm” package in Python for the implementation and trained the model using Bayesian optimization [48, 14] via the ”Optuna” package [1].
-For all methods, we conducted performance evaluation using 5-fold cross-validation.
+This prior provides flexibility in modeling the coefficient distribution and can be easily implemented using the Pyro probabilistic programming language [3].
+  • Logistic Regression with the Laplace (LP) prior: the performance of logistic regression was also explored using a standard Laplace prior distribution, which promotes sparsity in the coefficient estimates as a Bayesian Lasso [44].
+This prior has proven effective in high-dimensional settings and can be efficiently implemented using Pyro, leveraging its support for non-Gaussian prior distributions [3].
+  • LightGBM: As a baseline method for table data prediction tasks, LightGBM [28], which is an efficient and scalable gradient boosting framework, was implemented adopting the ”lightgbm” package in Python for the implementation and training of the model using Bayesian optimization [49, 15] via the ”Optuna” package [1].
+  • Principal Component Regression: As a representative method for high-dimensional data, PCR was implemented with the number of bases selected from $1$ to $30$ by 5-fold cross-validation.
+For all methods, performance evaluation was conducted using 5-fold cross-validation.
 In each fold, the model was trained on the training data, and predictions were made on the test data.
-As in the previous section, the evaluation metrics used to assess predictive performance were Accuracy, Area Under the ROC Curve (AUC), UM, and CC.
+As in the previous section, the evaluation metrics used to assess predictive performance were Accuracy, AUC, UM, and CC.
 6.3. Results
-The results presented in Table 1 provide valuable insights into the performance of the proposed Bayesian logistic regression approach with different prior distributions, as well as its comparison to the LightGBM baseline.
-Overall, Logistic Regression with the ES prior distribution achieves the highest accuracy and AUC among all the methods, demonstrating its effectiveness in capturing the underlying structure of the high-dimensional ARCENE dataset.
+The results of the experiment are presented in Table 1.
+The entries represent the average values of the metrics over $5$-fold cross-validation.
+Overall, among all the methods, logistic regression with the ES prior distribution achieves the highest accuracy and AUC, demonstrating its effectiveness in capturing the underlying structure of the high-dimensional ARCENE dataset.
 The superior performance of the ES prior can be attributed to its ability to adapt to the intrinsic dimensionality of the data by incorporating the empirical spectral information of the covariance matrix [60].
 Let us take a closer look at the comparison.
-The horseshoe prior, known for its flexibility and sparsity-inducing properties [12], also performs well, with the second-highest accuracy and AUC.
+The horseshoe prior, known for its flexibility and sparsity-inducing properties [13], performs well, with the second-highest accuracy and AUC.
 This suggests that the regression on ARCENE dataset may exhibit some degree of sparsity, which the HS prior is able to capture effectively.
@@
 Nevertheless, their performance is still superior to the LightGBM baseline, highlighting the advantages of the Bayesian approach in handling excess covariates.
+PCR effectively selects relevant components from a large number of covariates although the results are only moderate and do not match those of our method which retains all covariates without reduction.
 A notable aspect of the results is the uncertainty estimation of the Bayesian methods.
@@
 LightGBM 0.775 0.858
+PCR 0.810 0.878
 Table 1.
-Data Analysis: Data Analysis: Comparison of classification performance between our method, LightGBM.
-The metrics are Accuracy, Area Under the Curve (AUC), the proportion of uncertain predictions among misclassifications (UM), and the proportion of correct predictions among certain classifications (CC).
+Comparison of classification performance among logistic regressions with effective spectral (ES, our method), horseshoe (HS), Laplace (LP), and normal (NM) priors, LightGBM, and principal component regression (PCR), using accuracy, area under the curve (AUC), proportion of uncertain predictions among misclassifications (UM), and proportion of correct predictions among certain classifications (CC) as metrics.
 7. Discussion
-This study introduces a novel Bayesian approach for prediction in overparameterized generalized linear models, highlighting its theoretical foundations and practical applications.
+This study introduces a novel Bayesian approach for prediction in overparameterized generalized linear models, providing a theoretical foundation for practical applications.
 The primary theoretical contribution lies in establishing posterior contraction under appropriate conditions, validating the use of the posterior distribution as a reliable approximation of the true data-generating distribution.
 Empirical evidence from simulations and real-world datasets corroborates the effectiveness of the proposed method in delivering accurate predictions and reliable uncertainty estimates.
-In the realm of Bayesian high-dimensional inference, the proposed approach distinguishes itself by directly addressing non-sparsity while offering theoretical guarantees for prediction with nonlinearity.
-This unique perspective paves the way for new avenues of Bayesian inference in overparameterized models.
+In the field of high-dimensional Bayesian inference, the proposed approach stands out by directly addressing non-sparsity while providing theoretical guarantees for prediction under nonlinearity.
+This unique perspective paves the way for Bayesian inference in overparameterized models.
 This study marks a significant advance in overparameterization theory by achieving consistency in nonlinear regression problems.
-The robustness of posterior contraction theory to model complexity (nonlinearity) plays a pivotal role in attaining consistency because it solely necessitates control of the covering number, the decay of the prior distribution in its tail region, and the concentration of the prior distribution around the true value of the parameter.
-Finally, we conclude this paper by discussing its limitations and identifying potential areas for improvement.
+The robustness of posterior contraction theory to model complexity (nonlinearity) plays a pivotal role in attaining consistency, necessitating control of the covering number, decay of the prior distribution in the tail region, and concentration of the prior distribution around the true value of the parameter.
+Finally, the limitations of the paper can help identify potential areas for improvement.
 The theoretical analysis hinges on specific assumptions regarding the covariance matrix, which may not always hold in real-world scenarios.
-Future research could focus on relaxing these assumptions and extending the results to more general settings as deep neural networks generalize to a wide range of data.
+Future research can focus on relaxing these assumptions to extend the results to more general settings, as deep neural networks generalize to a wide range of data.
 Moreover, the computational complexity associated with sampling from the posterior distribution can pose challenges for large-scale datasets.
-Also, further improvements in uncertainty estimation could also be combined with the more flexible approximation such as normalizing flow [44, 45] and Stein Variational Gradient Descent [35].
+Further improvements in uncertainty estimation can also be combined with more flexible approximation approaches, such as normalizing flow [45, 46] and Stein variational gradient descent [36].
+Computer Programs
+The computer programs used in the paper for the numerical experiments in Section 5 and the application in Section 6 are publicly available on GitHub repository https://github.com/TomWaka/BA-Overparameterized-NonLinReg.
+The computational algorithm is outlined in the appendix.
 Acknowledgements
-I would like to express my sincere gratitude to my advisor, Masaaki Imaizumi, for his patient and rigorous supervision. This work was supported by JSPS KAKENHI (22J21090) and JST ACT-X (JPMJAX23CS).
+I would like to express my sincere gratitude to my advisor, Masaaki Imaizumi, for his patient and rigorous supervision.
+I am also grateful to Shonosuke Sugasawa for his helpful suggestions.
+This work was supported by JSPS KAKENHI (22J21090) and JST ACT-X (JPMJAX23CS).
 Appendix A Proof of the Main Results
-In this section, we provide proof of the main results presented in the previous sections.
+In this section, proofs are provided for the main results presented in the previous sections.
 The key is to establish the posterior contraction rate for the squared $L_{2}(P_{\bm{X}})$ norm of the difference between the true and estimated mean functions, i.e., $\mathbb{E}[\{g(\bm{X}^{\top}\widehat{\bm{\beta}})-g(\bm{X}^{\top}\bm{\beta}^{*})\}^{2}]$.
@@
 To this end, we invoke a general result on posterior contraction rates from [18].
-Let $\mathcal{P}=\mathbb{R}^{p}$ be the parameter space equipped with the distance $d_{\mathcal{P}}$ induced by the $\|\cdot\|_{\Sigma}$ norm, and define a sequence of restricted parameter spaces
+Let $\mathcal{P}=\mathbb{R}^{p}$ be the parameter space equipped with distance $d_{\mathcal{P}}$ induced by the $\|\cdot\|_{\Sigma}$ norm, further defining it as a sequence of restricted parameter spaces
 $\mathcal{P}_{n}:=\{\bm{h}\in\mathbb{R}^{p}:\|\bm{h}\|_{2}\leq H_{n}\},$ (23)
@@
 Recall from Assumption 1 that the empirical covariance matrix $\widehat{\Sigma}$ concentrates around $\Sigma$ with probability at least $1-\ell_{n}$, which controls the randomness of the posterior distribution.
-Theorem 2.1 of [17] states that the posterior contraction rate in Theorem 1 holds if there exists a sequence $H_{n}\to\infty$ such that the following three conditions are satisfied for any sufficiently large $n$ and some constant $C>0$:
+Theorem 2.1 of [18] states that the posterior contraction in Theorem 1 holds if there exists a sequence $H_{n}\to\infty$ such that the following three conditions are satisfied for any sufficiently large $n$ and some constant $C>0$:
 $\log\mathcal{N}(\varepsilon_{n},\mathcal{P}_{n},d_{\mathcal{P}})\leq n\varepsilon_{n}^{2},$ (24)
@@
 $\Pi_{\mathfrak{b}}\left(\left\{\bm{\beta}:\|\bm{\beta}-\bm{\beta}^{*}\|_{\Sigma}^{2}\leq\varepsilon_{n}^{2}\right\}\mid\widehat{\Sigma}\right)\geq\exp(-Cn\varepsilon_{n}^{2}).$ (27)
-The proof of (26) follows from a similar argument as in [59], which we outline below for completeness.
-The extension to the case where $\sigma^{2}$ is unknown is straightforward.
-If we take $[1/n,n]$ as $\mathcal{P}_{n}$ when we have an inverse Gaussian distribution as a prior distribution for $\sigma^{2}$, the posterior contraction holds because $\mathcal{P}_{n}$ dominates $\mathcal{P}$ as $n$ increases and the remaining part $\mathcal{P}\setminus\mathcal{P}_{n}$ has only the negligible mass of the prior.
+The proof of (27) follows from a similar argument as in [60].
+The extension to the case where $\sigma^{2}$ is unknown, is straightforward.
+Taking $[1/n,n]$ as $\mathcal{P}_{n}$ in the case of an inverse Gaussian distribution as a prior distribution for $\sigma^{2}$, the posterior contraction holds because $\mathcal{P}_{n}$ dominates $\mathcal{P}$ as $n$ increases and the remaining $\mathcal{P}\setminus\mathcal{P}_{n}$ part has only the negligible mass of the prior.
 Combining the above arguments, we conclude that the conditions of Theorem 2.1 in [18] are satisfied, and hence the posterior contraction in Theorem 1 holds.
-The explicit characterization of the contraction rate $\varepsilon_{n}$ in terms of the eigenvalues of $\Sigma$ follows from a careful analysis of the interplay between the rate $\rho_{n}$ in Assumption 1, the upper and lower bounds $(L_{\kappa},U_{\kappa})$ in Assumption 2, and the additional assumption on the eigenvalue decay stated in Theorem 1.
+The explicit characterization of the contraction rate $\varepsilon_{n}$ in terms of the eigenvalues of $\Sigma$ follows from a careful analysis of the interplay between the rate $\rho_{n}$ in Assumption 1, upper and lower bounds $(L_{\kappa},U_{\kappa})$ in Assumption 2, and the additional assumption on the eigenvalue decay stated in Theorem 1.
 Appendix B Consistency as a Distribution and Point Estimator
-Following the discussion on posterior distribution contraction, we further explore the convergence of distributions and estimators using various metrics.
+Following the discussion on posterior distribution contraction, the convergence of distributions and estimators is further explored using various metrics.
 To extend the discussion, let us revisit Example 1.
 Example 2.
-Recall that $K(P,P^{\prime})$ is the KL divergence, $V(P,P^{\prime})$ is the KL variation, and $H(P,P^{\prime})$ is the Hellinger distance between models with different parameters.
+Recall that $K(P,P^{\prime})$ is the KL divergence; $V(P,P^{\prime})$ is the KL variation; and $H(P,P^{\prime})$ is the Hellinger distance between models with different parameters.
   1. (1)
@@
   2. (2)
-    For binary regression with the logistic link, we have
+    For binary regression with a logistic link, we have
     $K(P,P^{\prime}),V(P,P^{\prime}),H(P,P^{\prime})\lesssim\|\bm{\beta}-\bm{\beta}^{\prime}\|_{\Sigma}^{2}.$
@@
 As shown in the previous section, the consistency of $\|\bm{\beta}-\bm{\beta}^{\prime}\|_{\Sigma}^{2}$ holds.
-Consequently, in the above examples, if the distance between distributions can be bounded by constant times the quantity, we obtain the consistency of the posterior distribution to the true data-generating distribution with respect to the KL divergence, KL variation, and Hellinger distance.
-Furthermore, because the total variation distance and the Hellinger distance induce the same topology, posterior consistency also holds in terms of the total variation distance.
+Consequently, in the above examples, if the distance between distributions can be bounded by a constant times the quantity, the consistency of the posterior distribution to the true data-generating distribution is obtained with respect to the KL divergence, KL variation, and Hellinger distance.
+Furthermore, because the total variation and Hellinger distances induce the same topology, posterior consistency also holds in terms of the total variation distance.
 In connection with distributional convergence, we now shift our focus to the convergence of the posterior mean.
-If the parameter space $\mathcal{P}$ is convex and the distance $d_{\mathcal{P}}$ on $\mathcal{P}$ is bounded and convex, then the posterior mean serves as a consistent estimator ([e.g., Theorem 6.8 in 20]).
-The assumption of $\mathcal{P}$ being convex is often satisfied, and the boundedness and convexity of the distance hold, for instance, with the total variation distance.
+If the parameter space $\mathcal{P}$ is convex and distance $d_{\mathcal{P}}$ on $\mathcal{P}$ is bounded and convex, then the posterior mean serves as a consistent estimator ([e.g., Theorem 6.8 in 21]).
+The assumption of $\mathcal{P}$ being convex is often satisfied, and the boundedness and convexity of the distance hold for the total variation distance.
 Hence, in the above examples, the consistency of the posterior mean is achieved.
 Appendix C Computational Method
-In this section, we describe the computational method for the proposed approach.
-We employ stochastic variational inference [19, 21, 4] to approximate the posterior distribution of the model parameters.
-Since the closed-form expression of the KL divergence is not available, we use the reparameterization trick [28, 46] to enable efficient gradient-based optimization.
-The reparameterization trick allows us to express the variational distribution in terms of a deterministic function of the variational parameters and a random variable, making the gradients of the variational parameters computable.
-The optimization part is performed using the Adam algorithm [25].
+In this section, the computational method for the proposed approach is described.
+Stochastic variational inference [20, 22, 4] is employed to approximate the posterior distribution of the model parameters.
+Since the closed-form expression of the KL divergence is not available, a reparameterization trick [29, 47] was used to enable efficient gradient-based optimization.
+The reparameterization trick allows the variational distribution to be expressed in terms of a deterministic function of the variational parameters and a random variable, making the gradients of the variational parameters computable.
+The optimization part was performed using the Adam optimizer [26].
 C.1. Logistic Regression Example
-To illustrate the computational procedure, we consider the example of logistic regression. The model components are as follows:
+To illustrate the computational procedure, let us consider an example of logistic regression. The model components are as follows:
   1. (1)
@@
 However, it can be approximated as a normal distribution if $R$ is taken large enough.
-Also, a procedure to compute the posterior distribution without support constraints and then project it onto the constrained region is justified by [49].
-Our goal is to find the optimal variational parameters ${\widetilde{\bm{\mu}}}_{opt}$ and $\widetilde{\Sigma}_{opt}$ that minimize the KL divergence between the variational approximation $q(\bm{\beta}\mid{\widetilde{\bm{\mu}}},\widetilde{\Sigma})$ and the true posterior $p_{pos}$:
+Also, a procedure that computes the posterior distribution without support constraints and then projects it onto the constrained region, was justified by [50].
+Our objectove is to find the optimal variational parameters ${\widetilde{\bm{\mu}}}_{opt}$ and $\widetilde{\Sigma}_{opt}$ that would minimize the KL divergence between the variational approximation $q(\bm{\beta}\mid{\widetilde{\bm{\mu}}},\widetilde{\Sigma})$ and the true posterior $p_{pos}$:
 ${\widetilde{\bm{\mu}}}_{opt},\widetilde{\Sigma}_{opt}:=\operatornamewithlimits{argmin}_{{\widetilde{\bm{\mu}}},\widetilde{\Sigma}}\mathrm{KL}(q(\bm{\beta}\mid{\widetilde{\bm{\mu}}},\widetilde{\Sigma})\|p_{pos}),$ (28)
@@
 $\mathrm{KL}(q(\bm{\beta}\mid{\widetilde{\bm{\mu}}},\widetilde{\Sigma})\|p_{pos})=\mathbb{E}_{q}[\log q]-\mathbb{E}_{q}[\log\phi(\bm{\beta}\mid\bm{0},\widehat{\Sigma}_{1:k})]-\sum_{i=1}^{n}\mathbb{E}_{q}[\log L(y_{i};\bm{x}_{i}^{\top}\bm{\beta})]+\mathrm{const}.$ (29)
-To compute the gradients of the KL divergence with respect to the variational parameters, we employ the reparameterization trick.
-Specifically, we denote by $\widetilde{\bm{\beta}}$ a sample from the variational distribution $q$ and parameterize it as $\widetilde{\bm{\beta}}=\widetilde{\bm{\mu}}+\widetilde{\Sigma}_{1/2}\widetilde{\bm{\epsilon}}$, where $\widetilde{\bm{\epsilon}}$ is a sample from a standard Gaussian distribution $\mathcal{N}(\bm{0},\bm{I}_{p})$ and $\widetilde{\Sigma}_{1/2}$ is a decomposition of $\widetilde{\Sigma}$, i.e., $\widetilde{\Sigma}=\widetilde{\Sigma}_{1/2}\widetilde{\Sigma}_{1/2}^{\top}$.
+To compute the gradients of the KL divergence with respect to the variational parameters, we employ a reparameterization trick.
+Specifically, denoting by $\widetilde{\bm{\beta}}$ a sample from the variational distribution $q$, we parameterize it as $\widetilde{\bm{\beta}}=\widetilde{\bm{\mu}}+\widetilde{\Sigma}_{1/2}\widetilde{\bm{\epsilon}}$, where $\widetilde{\bm{\epsilon}}$ is a sample from a standard Gaussian distribution $\mathcal{N}(\bm{0};\bm{I}_{p})$ and $\widetilde{\Sigma}_{1/2}$ are obtained from a decomposition of $\widetilde{\Sigma}$, i.e., $\widetilde{\Sigma}=\widetilde{\Sigma}_{1/2}\widetilde{\Sigma}_{1/2}^{\top}$.
 With this reparameterization, the KL divergence can be approximated as an unbiased estimate
 $\mathrm{KL}(q(\bm{\beta}\mid{\widetilde{\bm{\mu}}},\widetilde{\Sigma})\|p_{pos})\approx\log q(\widetilde{\bm{\beta}}\mid{\widetilde{\bm{\mu}}},\widetilde{\Sigma})-\log\phi(\widetilde{\bm{\beta}}\mid\bm{0},\widehat{\Sigma}_{1:k})-\sum_{i=1}^{n}\log L(y_{i};\bm{x}_{i}^{\top}\widetilde{\bm{\beta}})+\mathrm{const}.$ (30)
-We can now compute the gradients of the KL divergence with respect to the variational parameters. The gradient with respect to $\widetilde{\bm{\mu}}$ is given by
+Now, the gradients of the KL divergence can be computed with respect to the variational parameters. The gradient with respect to $\widetilde{\bm{\mu}}$ is given by
 $\frac{\partial}{\partial\widetilde{\bm{\mu}}}\mathrm{KL}(q(\widetilde{\bm{\beta}}\mid{\widetilde{\bm{\mu}}},\widetilde{\Sigma})\|p_{pos})$ (31)
@@
 The derived gradients can be used in conjunction with the Adam optimizer to update the variational parameters $\widetilde{\bm{\mu}}$ and $\widetilde{\Sigma}_{1/2}$.
-The optimization procedure iteratively updates the variational parameters until convergence or a maximum number of iterations is reached.
+The optimization procedure iteratively updates the variational parameters until the maximum number of iterations or convergence is reached.